### v1

# Comparing Bayesian and Frequentist Inference in Biological Models: A Comparative Analysis of Accuracy, Uncertainty, and Identifiability

Abstract
Mathematical models inform inference and forecasting in ecology and epidemiology, but results depend on the estimation framework. Here, we compare Bayesian and Frequentist approaches across three biological models using four datasets: Lotka–Volterra predator–prey dynamics (Hudson Bay), a generalized logistic model (lung injury and 2022 U.S. mpox), and an SEIUR epidemic model (COVID-19 in Spain). Both approaches use a normal error structure to ensure a fair comparison. We first assessed structural identifiability to determine which parameters can theoretically be recovered from the data. We then evaluated practical identifiability and forecasting performance using four metrics: mean absolute error (MAE), mean squared error (MSE), 95% prediction interval (PI) coverage, and weighted interval score (WIS). For the Lotka–Volterra model with both prey and predator data, we analyzed three observation scenarios: prey only, predator only, and both simultaneously. The Frequentist workflow was implemented using QuantDiffForecast (QDF) in MATLAB, which fits ODE models via nonlinear least squares and quantifies uncertainty through parametric bootstrap. The Bayesian workflow was implemented using BayesianFitForecast (BFF), which employs Hamiltonian Monte Carlo sampling via Stan to generate posterior distributions and diagnostics such as the Gelman–Rubin $\hat{R}$ statistic. Results show that Frequentist inference performs best in well-observed settings with rich data (e.g., GLM for lung injury and mpox, or Lotka–Volterra when both species are observed). In contrast, Bayesian inference excels when latent-state uncertainty is high and data are sparse or partially observed, as seen in the SEIUR model for COVID-19. Structural identifiability analysis clarifies these patterns: full observability enhances both frameworks, while limited observability constrains parameter recovery regardless of the method. This comparative analysis provides practical guidance for selecting inference frameworks based on data richness, observability, and uncertainty quantification needs.
1 Introduction
Mathematical models based on ordinary differential equations (ODEs) are essential tools for understanding and forecasting dynamics in ecology and epidemiology (modeling_overview; mohammed2024trophiccascade; alsammani2025cholera). Model-based forecasts have been instrumental in managing major public health crises over the past decade. During the COVID-19 pandemic, forecasts guided resource allocation and social distancing policies (dixon2022comparison; cheng2023real; lutz2019applying; shearer2020infectious; chowell2022ensemble; chowell2022sub; alsammani2025cholera). The U.S. CDC’s FluSight Challenge used models to optimize influenza vaccine distribution and public health messaging (reich2019collaborative; mcgowan2019collaborative; biggerstaff2014estimates). During the West African and DRC Ebola outbreaks, models predicted transmission patterns and evaluated intervention effectiveness (chowell2017perspectives; funk2019assessing; meltzer2014estimating; chretien2015mathematical; chowell2014west; roosa2020multi). More recently, forecasting models for mpox predicted its spread and assessed containment measures (bleichrodt2023real; bleichrodt2023retrospective; charniga2024nowcasting; chowell2024growthpredict). In ecology, predator-prey models help predict population dynamics and inform conservation strategies (lotka_volterra_data). These applications demonstrate that reliable forecasts depend on accurate parameter estimation (forecast_reliability; Chowell2024QuantDiffForecast; Karami2025BayesianFitForecast).
However, the ability to recover meaningful parameters from data critically depends on identifiability—the question of whether model parameters can be uniquely determined from available observations (identifiability_review; practical_identifiability; atanasov2025honeybee). Identifiability can fail when data are sparse, models are over-parameterized, or parameters are strongly correlated, leading to misleading inferences and unreliable forecasts. Recent work has shown that factors such as vaccination behavior, immunity duration, and data completeness shape both epidemic trajectories and the reliability of parameter estimates (Alsammani2025MBE). Understanding how identifiability interacts with estimation frameworks is essential to ensure that model-based forecasts are interpretable and reproducible. Structural or practical identifiability limitations can cause two inference frameworks to produce different results even under identical models and data.
Two estimation paradigms dominate the field: Bayesian and Frequentist methods. Frequentist methods typically calibrate ODE models by optimizing a likelihood function or minimizing an objective function such as the sum of squared differences between observed and predicted values (gneiting2008probabilistic; mwambi2011frequentist; chowell2017fitting; chowell2020real). These methods use algorithms like gradient descent or the Levenberg-Marquardt algorithm, assume specific distributions for measurement errors (e.g., Gaussian or Poisson), and quantify uncertainty through bootstrapping techniques (banks2014modeling; pruitt2024role; transtrum2012optimal; huang2024nonlinear; bates1988nonlinear; cao2012penalized; ramsay2017dynamic; seber2003nonlinear). Frequentist methods are computationally efficient and often perform well when data are abundant and high quality (deterministic_text; Chowell2024QuantDiffForecast). The QuantDiffForecast (QDF) toolbox implements this workflow for fitting models and generating predictions with quantified uncertainty (chowell2024parameter).
Bayesian methods, in contrast, apply Bayes’ theorem to combine prior distributions of parameters with the likelihood of observed data, producing posterior distributions that explicitly incorporate uncertainty (greenland2009bayesian; mckinley2014simulation; kypraios2017tutorial; girolami2008bayesian; grinsztajn2021bayesian; bouman2024bayesian; gelman2020bayesian; belasso2023bayesian). These methods typically use Markov Chain Monte Carlo (MCMC) algorithms to approximate posterior distributions, providing comprehensive measures of parameter uncertainty and credible intervals (martin2020computing; dunson2001commentary; harel2018multiple; grinsztajn2006bayesian; annis2017bayesian; kelter2020bayesian). Bayesian methods can better navigate complex parameter spaces, avoid local optima, and handle incomplete or noisy data (gelman1996physiological; huang2006hierarchical; huang2020bayesian). Tools like Stan facilitate Bayesian estimation and forecasting, allowing rigorous uncertainty quantification and model validation (bayesian_text; gelman_rhat; posterior_diagnostics; Karami2025BayesianFitForecast; monnahan2017faster; burkner2017brms). The BayesianFitForecast (BFF) toolbox implements this workflow with diagnostics such as the Gelman-Rubin $\hat{R}$ statistic.
Despite extensive use of both paradigms, prior comparisons often vary models, likelihoods, or preprocessing between methods, making it difficult to attribute observed differences to the estimation framework itself (benchmarking_pitfalls). To address this gap, we conduct a controlled comparison of Bayesian and Frequentist inference under standardized conditions: the same models, the same normal error structure, and harmonized data preprocessing. We analyze three systems and four datasets representing different levels of complexity and observability: the Lotka-Volterra (LV) predator-prey model (Hudson Bay lynx-hare data), a generalized logistic model (GLM) for lung injury and the 2022 U.S. mpox outbreak, and an SEIUR epidemic model for the first COVID-19 wave in Spain (lotka_volterra_data; lung_injury_glm; mpox_us_2022; spain_covid_firstwave). For the LV model, we analyze three observation scenarios (prey only, predator only, and both simultaneously) to assess how partial observability affects parameter recovery.
We integrate structural identifiability analysis to determine which parameters can theoretically be recovered from the data, separating fundamental data limits from algorithmic limits (identifiability_methods; stigter2023structural; dankwa2022identifiability). We then evaluate practical identifiability and forecasting performance using four metrics: mean absolute error (MAE), mean squared error (MSE), 95% prediction interval (PI) coverage, and weighted interval score (WIS) (wis_ref; coverage_ref; mcmc_convergence; Chowell2024QuantDiffForecast; Karami2025BayesianFitForecast; Gneiting2007).
Our objectives are to: (i) compare prediction accuracy across Bayesian and Frequentist inference, (ii) assess uncertainty calibration and diagnostics, and (iii) relate observed performance to structural versus practical identifiability under full versus partial observation. This provides practical guidance on when each paradigm is preferable based on data richness, observability, and uncertainty quantification needs (Karami2025BayesianFitForecast; Chowell2024QuantDiffForecast).
The remainder of this paper is organized as follows. Section 2 describes the four datasets analyzed in this study, including the Hudson Bay lynx-hare data, lung injury outbreak data, mpox epidemic data, and COVID-19 data from Spain. Section 3 presents the three mathematical models employed: the Lotka–Volterra predator-prey model, the generalized logistic model, and the SEIUR epidemic model. Section 4 details the Bayesian and Frequentist inference methodologies, including likelihood specifications, prior distributions, uncertainty quantification procedures, and performance metrics. Section 5 presents the structural identifiability analysis for all models under different observation scenarios, establishing theoretical bounds on parameter recovery. Section 6 reports empirical results, including parameter estimates, performance metrics, and convergence diagnostics for all models and datasets. In Section 7, the Discussion synthesizes our findings, interprets the role of structural identifiability in explaining performance differences, and provides practical guidance for method selection. The Conclusion, Section 8, summarizes the main contributions and implications for biological modeling practice.
2 Data
We analyzed four datasets spanning ecological and epidemiological systems, each selected to represent different levels of data richness, temporal resolution, and observability (Table 1; Figure 1). All datasets consist of time-series observations at discrete time points $t_{n}$, where $n=1,\ldots,N$ denotes the observation index. The datasets include both population counts (for the ecological system) and incident case counts (for epidemic systems). Data preprocessing was standardized across all datasets to ensure comparability between inference methods; specific preprocessing steps are summarized at the end of this section.
2.1 Hudson Bay Lynx–Hare Data
The Hudson Bay lynx–hare dataset consists of annual counts of Canadian lynx (Lynx canadensis) and snowshoe hare (Lepus americanus) populations from 1900 to 1920, derived from Hudson’s Bay Company pelt return records (carpenter2018predator; elton1924periodic). These records serve as a proxy for population abundance, with pelt counts reflecting relative population size. The dataset contains $N=21$ yearly observations for both prey (hare) and predator (lynx) populations. This dataset is particularly valuable for assessing parameter identifiability in predator–prey dynamics because it provides simultaneous observations of both interacting species over multiple population cycles. The data have been widely used as a benchmark for testing ecological models and parameter estimation methods. No smoothing, interpolation, or imputation was applied to the original data.
2.2 Lung Injury Data (EVALI)
This dataset comprises weekly incident case counts of electronic-cigarette or vaping product use-associated lung injury (EVALI) reported in the United States during 2019 (krishnasamy2020evali; lozier2019evaliupdate; chowell2020evalimodel). The outbreak was first identified in mid-June 2019, with cases rapidly escalating through the summer and fall before declining by November 2019. We analyzed $N\approx 21$ weeks of data spanning mid-June to early November 2019, as reported through the Centers for Disease Control and Prevention (CDC) Morbidity and Mortality Weekly Report (MMWR) surveillance system. Case definitions followed CDC criteria for confirmed and probable EVALI cases. The data were aggregated by epidemiological week (MMWR week) with no backfill corrections applied beyond those included in the official CDC reports. This dataset represents a short-duration outbreak with a clear peak, making it suitable for testing phenomenological growth models under well-defined epidemic dynamics.
2.3 Mpox Data (United States, 2022)
We analyzed weekly incident confirmed or probable mpox cases in the United States during the 2022–2023 outbreak as reported in (cdc2022mpox; chiu2023mpox). The outbreak began in May 2022 and peaked in August 2022 before declining through early 2023. Data were compiled from CDC national surveillance reports aggregated by MMWR week. Case definitions followed CDC guidelines for confirmed (laboratory-confirmed orthopoxvirus with specimen typed as monkeypox virus) and probable cases (epidemiologically linked to a confirmed case without laboratory confirmation). The temporal window analyzed corresponds to the public reporting period documented in CDC surveillance dashboards (see references for access dates). This dataset provides an example of an emerging infectious disease outbreak with intensive surveillance and public health response, allowing evaluation of model performance during rapidly evolving epidemic conditions.
2.4 COVID-19 Data (Spain, First Wave)
This dataset consists of daily incident laboratory-confirmed COVID-19 cases reported during the first epidemic wave in Spain from February through May 2020 (espana2020covid; chowell2020covidspain). Data were compiled from official reports by the Spanish Ministry of Health (Ministerio de Sanidad), with case dates based on report date rather than symptom onset date due to data availability constraints during the early pandemic response. The first wave in Spain exhibited rapid exponential growth followed by a peak in late March 2020 and subsequent decline following implementation of strict non-pharmaceutical interventions including a national lockdown. We restricted analysis to the first wave period to avoid complications from changing surveillance protocols, intervention policies, and emergence of new variants that occurred in later waves. Known reporting artifacts explicitly flagged in official data releases were excluded. This dataset represents a large-scale epidemic with significant underreporting and latent compartments (exposed and unreported infectious individuals), making it ideal for assessing inference methods under partial observability and model complexity.
Dataset Source & Period Frequency Observations (N)
Hudson Bay Lynx–Hare Hudson’s Bay Company pelt returns (1900–1920) Annual 21
Lung Injury (EVALI, US) CDC MMWR reports (mid-Jun–Nov 2019) Weekly $\approx$21
Mpox Epidemic (US, 2022) CDC surveillance (May 2022–early 2023) Weekly Variable
COVID-19 (Spain, 1st Wave) Spanish Ministry of Health (Feb–May 2020) Daily Variable
Table 1: Summary of datasets analyzed in this study, including data source, temporal coverage, reporting frequency, and number of observations for each system.
2.5 Preprocessing Summary
All datasets were processed using a standardized protocol to ensure comparability across inference methods. The preprocessing steps were as follows:
2.5.0.1 Temporal aggregation
The Hudson Bay lynx–hare data were used as published without aggregation. Lung injury and mpox data were aggregated at the weekly level (MMWR epidemiological week) as reported in official CDC surveillance. COVID-19 data for Spain were analyzed at the daily level as reported by the Spanish Ministry of Health.
2.5.0.2 Data type
The Hudson Bay dataset represents population abundance (pelt counts), while all epidemic datasets (lung injury, mpox, COVID-19) represent incident case counts per reporting period. For the generalized logistic model (GLM), cumulative case counts were derived by summing incident counts, and model predictions were compared to incident data via the time derivative of cumulative cases.
2.5.0.3 Missing values
No missing values were present in the final temporal windows used for model fitting. No imputation was performed.
2.5.0.4 Smoothing
No smoothing was applied to the data used for parameter estimation or forecasting. Figures may display smoothed overlays for visualization purposes only; these smoothed curves were not used in model fitting.
2.5.0.5 Temporal windows
Modeling windows for calibration and forecasting were selected based on epidemic phase and data availability. Specific windows for each dataset are described in the Results section.
2.5.0.6 Data quality control
For COVID-19 data, known reporting artifacts explicitly flagged in official Spanish Ministry of Health releases were excluded. For all other datasets, data were used as published in official surveillance reports without additional corrections.
Figure 1: Time series of ecological and epidemic population dynamics. The Hudson Bay Lynx–Hare dataset shows annual prey and predator abundance (1900–1920), while the remaining panels depict weekly or daily reported cases from major disease outbreaks: Lung Injury (Electronic cigarette or vaping product use-associated lung injury (EVALI), US, 2019), Mpox (US, 2022), and COVID-19 (Spain, 2020).
3 Models
In this study, we employed three compartmental biological and epidemiological models to comprehensively assess the parameter identifiability of our two approaches: (i) the LV model, describing predator-prey interactions and population feedback mechanisms; (ii) the Generalized Logistic Model, capturing flexible epidemic growth through nonlinear case dynamics; and (iii) the SEIUR model, explicitly tracking susceptible, exposed, infectious (reported and unreported), and recovered populations to account for both observed and hidden transmission. These models were selected to represent a progression in model complexity and data observability: from a fully observed ecological system (LV), to a parsimonious single-equation epidemic model (GLM), to a multi-compartment latent-state model with partial observability (SEIUR). This progression enables a systematic comparison of Bayesian and Frequentist inference performance across increasing levels of model structure and latent uncertainty.
3.0.0.1 LV model
The LV equations (Lotka 1925; Volterra 1926, 1927) consist of a pair of first-order ODE describing the population dynamics of two interacting species: one predator and one prey. Let $x$ denote the prey population and $y$ the predator population at time $t$. Volterra modeled the temporal dynamics of these populations as follows:
$\frac{dx}{dt}=\alpha x-\beta xy,\quad\frac{dy}{dt}=-\gamma y+\delta xy,$ (1)
with initial conditions
$x(0)=x_{0},\quad y(0)=y_{0},$ (2)
where $\alpha$ is the intrinsic growth rate of the prey population, $\beta$ is the predation rate coefficient, $\gamma$ is the mortality rate of the predator population, and $\delta$ is the predator’s growth rate per prey consumed. The initial conditions $x_{0}$ and $y_{0}$ specify the starting population sizes of prey and predator, respectively. The LV model serves as a classical benchmark in population ecology, offering a well-understood test case for examining parameter identifiability and uncertainty propagation under full versus partial observability.
xxyyIntrinsic growth $\alpha x$Loss due to predation $\beta xy$Growth from consuming prey $\delta xy$Natural mortality $\gamma y$
Figure 2: LV predator–prey diagram. Circles represent the prey $x$ and predator $y$ populations. The self-loop on $x$ indicates intrinsic growth at rate $\alpha$. Curved arrows between $x$ and $y$ represent interactions: prey loss due to predation and predator growth from consuming prey. The downward arrow on $y$ represents natural mortality.
3.0.0.2 Generalized Logistic Model
The Generalized Logistic Model (GLM) is a flexible extension of the logistic function used to model S-shaped epidemic growth curves. It is given by
$\frac{dC}{dt}=r\,C^{p}(t)\left(1-\frac{C(t)}{K}\right),$ (3)
with initial condition
$C(0)=C_{0},$ (4)
where $C(t)$ is the cumulative number of cases at time $t$, $r$ is the generalized growth rate, $K$ is the final epidemic size, and $p\in[0,1]$ controls the growth dynamics: $p=0$ corresponds to constant incidence, $0<p<1$ to sub-exponential growth, and $p=1$ to exponential growth. Here, $C_{0}$ represents the first reported case. Moreover, $\frac{dC}{dt}$ serves as the observation operator mapping the model states to the reported incidence data. The GLM was chosen because it provides a parsimonious, single-equation description of epidemic trajectories and serves as a tractable case for testing inference accuracy and uncertainty when the full time series is directly observed. This model has also been widely applied in epidemiological settings to capture early and full-phase epidemic dynamics chowell2016zika . The GLM has been widely used as a phenomenological growth model to capture epidemic curves with rapid increase and saturation chowell2016zika; pell2018ebola .
[…]
Figure 3: Diagram of the GLM. The circle represents the cumulative cases $C(t)$. The self-loop indicates growth governed by the generalized logistic equation. The dashed arrow indicates the observed incidence.
3.0.0.3 SEIUR Model
The SEIUR model tracks susceptible ($S$), exposed ($E$), reported infectious ($I$), unreported infectious ($U$), recovered ($R$), and cumulative reported cases ($C$). The dynamics are
$\begin{gathered}\frac{dS}{dt}=-\beta_{f}(t)\frac{(I+U)S}{N},\quad\frac{dE}{dt}=\beta_{f}(t)\frac{(I+U)S}{N}-\kappa E,\quad\frac{dI}{dt}=\kappa\rho E-\gamma I,\\ \frac{dU}{dt}=\kappa(1-\rho)E-\gamma U,\quad\frac{dR}{dt}=\gamma(I+U),\quad\frac{dC}{dt}=\kappa\rho E,\end{gathered}$ (5)
with initial conditions
$(S(0),E(0),I(0),U(0),R(0),C(0))=(N-C_{0},0,C_{0},0,0,C_{0}),$ (6)
where $N$ is the total population, $\beta_{f}(t)$ is the time-dependent transmission rate, $\kappa$ is the incubation rate, $\gamma$ is the recovery rate, $\rho$ is the reporting proportion, and $C_{0}$ is the first reported case. The initial condition assumes that all initially infected individuals are reported, and the remaining population is susceptible. The time-dependent transmission rate $\beta_{f}(t)$ is defined as the following
$\beta_{f}(t)=\begin{cases}\beta_{0},&\text{if }t<t_{\text{int}},\\[6.0pt] \beta_{1}+(\beta_{0}-\beta_{1})\,e^{-q_{1}(t-t_{\text{int}})},&\text{if }t\geq t_{\text{int}}.\end{cases}$
This model represents a more realistic epidemic process with both observed and hidden states, making it ideal for testing how Bayesian and Frequentist methods handle latent variables, parameter coupling, and partial identifiability.
[…]
Figure 4: Compartmental diagram of the SEIUR model with underreporting. Circles represent the epidemiological compartments. Solid arrows indicate transitions between compartments, and the dashed arrow indicates the source of observed cases.
4 Methods
This section describes the Bayesian and Frequentist estimation frameworks used for parameter inference, uncertainty quantification, and forecasting. Let $Y=(y_{t_{1}},\ldots,y_{t_{n}})$ denote the observed data at discrete time points $t_{1},\ldots,t_{n}$, and let $\boldsymbol{\theta}$ represent the vector of unknown model parameters to be estimated. For the Lotka–Volterra (LV) model, $\boldsymbol{\theta}=(\alpha,\beta,\gamma,\delta)$. For the generalized logistic model (GLM), $\boldsymbol{\theta}=(r,p,K)$. For the SEIUR model, $\boldsymbol{\theta}=(\beta_{0},\beta_{1},q_{1},\rho,\kappa,\gamma_{1})$. Both estimation approaches assume a normal (Gaussian) error structure to ensure comparability of results.
4.1 Bayesian Inference
Bayesian inference integrates prior knowledge with observed data to obtain the posterior distribution of model parameters vanDeSchoot2021. This probabilistic framework is particularly effective when prior information is available or when data are sparse, noisy, or subject to partial observability, as it provides comprehensive uncertainty quantification through posterior distributions gelman2020bayesian; grinsztajn2021bayesian.
4.1.1 Bayes’ Rule
According to Bayes’ theorem, the posterior distribution of parameters given the data is proportional to the product of the prior distribution and the likelihood:
$p(\boldsymbol{\theta}\mid Y)\propto p(\boldsymbol{\theta})\,p(Y\mid\boldsymbol{\theta}),$ (7)
where $p(\boldsymbol{\theta})$ is the prior distribution encoding existing knowledge about the parameters before observing data, $p(Y\mid\boldsymbol{\theta})$ is the likelihood function representing the probability of observing the data given parameter values, and $p(\boldsymbol{\theta}\mid Y)$ is the posterior distribution of parameters after incorporating the observed data.
4.1.2 Likelihood Function
We assume that observation errors are independent and identically distributed (i.i.d.) following a normal distribution with constant variance. For the LV model with observed prey and predator populations, the likelihood is:
$y_{t_{j}}^{\text{prey}}\mid\boldsymbol{\theta}\sim\mathcal{N}(x_{\boldsymbol{\theta}}(t_{j}),\sigma_{\text{prey}}^{2}),\quad y_{t_{j}}^{\text{predator}}\mid\boldsymbol{\theta}\sim\mathcal{N}(y_{\boldsymbol{\theta}}(t_{j}),\sigma_{\text{predator}}^{2}),$ (8)
independently for $j=1,\ldots,n$, where $x_{\boldsymbol{\theta}}(t)$ and $y_{\boldsymbol{\theta}}(t)$ denote the prey and predator population solutions to the LV ODE system given parameters $\boldsymbol{\theta}$, and $\sigma_{\text{prey}}^{2}$ and $\sigma_{\text{predator}}^{2}$ are observation error variances.
For the GLM, the likelihood is based on incident cases (the time derivative of cumulative cases):
$y_{t_{j}}\mid\boldsymbol{\theta}\sim\mathcal{N}\left(\frac{dC_{\boldsymbol{\theta}}(t_{j})}{dt},\sigma^{2}\right),\quad j=1,\ldots,n,$ (9)
where $C_{\boldsymbol{\theta}}(t)$ is the solution to the GLM ODE given parameters $\boldsymbol{\theta}$.
For the SEIUR model, the likelihood is based on incident reported cases:
$y_{t_{j}}\mid\boldsymbol{\theta}\sim\mathcal{N}\left(\kappa\rho E_{\boldsymbol{\theta}}(t_{j}),\sigma^{2}\right),\quad j=1,\ldots,n,$ (10)
where $E_{\boldsymbol{\theta}}(t)$ is the exposed compartment solution to the SEIUR ODE system given parameters $\boldsymbol{\theta}$, and $\kappa\rho E$ represents the flow of individuals from exposed to reported infectious status.
4.1.3 Prior Distributions
Prior distributions encode existing knowledge or beliefs about parameter values before observing data greenland2009bias . For parameters with established biological or epidemiological interpretations, we specify informative priors based on previous studies. When prior knowledge is limited, we use weakly informative or uniform priors over plausible parameter ranges. For structural identifiability analysis with simulated data, priors were centered on true parameter values with varying degrees of precision to assess the impact of prior strength on parameter recovery. Table 14,18,19,23,24,28,29 summarizes the parameter bounds used for each model. The normal error variance $\sigma^{2}$ is assigned a weakly informative inverse-gamma or half-Cauchy prior to allow the data to inform the observation error scale.
4.1.4 Posterior Sampling and Convergence Diagnostics
Posterior distributions are approximated using Markov Chain Monte Carlo (MCMC) sampling implemented through the Hamiltonian Monte Carlo (HMC) algorithm in Stan vehtari2021rhat; burkner2017brms. For each model, we ran four independent chains with a sufficient number of iterations (typically 2,000–4,000) after a warm-up (burn-in) period to ensure convergence to the stationary distribution. Convergence was assessed using the Gelman–Rubin diagnostic $\hat{R}$ statistic burkner2017brms, which compares between-chain and within-chain variance; values of $\hat{R}<1.05$ (ideally $\hat{R}\approx 1.01$) indicate successful convergence. Effective sample size (ESS) was monitored to ensure adequate sampling of the posterior distribution. Trace plots and posterior density plots were visually inspected to confirm mixing and convergence.
Posterior summaries are reported as medians with 95% credible intervals (CrIs), which represent the central 95% of the posterior distribution. For forecasting, we propagate parameter uncertainty by sampling parameters values from the posterior distribution, solving the ODE for each sample, and summarizing the resulting forecast distribution with median and 95% prediction intervals (PIs).
4.1.5 Computational Implementation
Bayesian inference was conducted using the (BFF) toolbox Karami2025BayesianFitForecast, an R package designed for fitting and forecasting ODE-based epidemic models. The BFF toolbox provides an automated workflow that generates Stan code based on user-specified model structures, priors, and data inputs, eliminating the need for users to program directly in Stan. The toolbox outputs include posterior parameter distributions, convergence diagnostics ($\hat{R}$, ESS), trace plots, posterior density plots, and forecasting results with quantified uncertainty. Performance metrics including mean absolute error (MAE), mean squared error (MSE), weighted interval score (WIS), and 95% PI coverage are automatically computed to evaluate model fit and forecast accuracy.
4.2 Frequentist Inference
Frequentist estimation treats model parameters as fixed but unknown quantities and estimates them by optimizing an objective function based solely on the observed data, without incorporating prior information chowell2017primer; chowell2020ensembles . Uncertainty quantification is performed through bootstrap resampling procedures.
4.2.1 Parameter Estimation
Under the assumption of normally distributed observation errors with constant variance (consistent with the Bayesian approach), parameter estimates are obtained by minimizing the sum of squared residuals between observed data and model predictions. This corresponds to nonlinear least squares (NLS) estimation:
$\hat{\boldsymbol{\theta}}=\underset{\boldsymbol{\theta}}{\text{argmin}}\sum_{j=1}^{n}\left(y_{t_{j}}-\mu_{j}(\boldsymbol{\theta})\right)^{2},$ (11)
where $\mu_{j}(\boldsymbol{\theta})$ is the model-predicted value at time $t_{j}$ given parameters $\boldsymbol{\theta}$. For the LV model, $\mu_{j}(\boldsymbol{\theta})$ consists of predicted prey and predator populations $(x_{\boldsymbol{\theta}}(t_{j}),y_{\boldsymbol{\theta}}(t_{j}))$. For the GLM, $\mu_{j}(\boldsymbol{\theta})=dC_{\boldsymbol{\theta}}(t_{j})/dt$. For the SEIUR model, $\mu_{j}(\boldsymbol{\theta})=\kappa\rho E_{\boldsymbol{\theta}}(t_{j})$.
The optimization is performed using gradient-based algorithms such as the Levenberg–Marquardt algorithm or trust-region methods, which iteratively search the parameter space to minimize the objective function. The observed data variance is estimated as:
$\hat{\sigma}^{2}=\frac{1}{n}\sum_{j=1}^{n}\left(y_{t_{j}}-\mu_{j}(\hat{\boldsymbol{\theta}})\right)^{2}.$ (12)
4.2.2 Uncertainty Quantification via Parametric Bootstrap
To quantify parameter uncertainty and construct confidence intervals, we employ a parametric bootstrap procedure chowell2017primer; banks2014uncertainty. This approach generates synthetic datasets by resampling from the fitted model and re-estimates parameters for each synthetic dataset. The distribution of bootstrap parameter estimates characterizes sampling variability. The steps are as follows:
4.2.2.1 Generate bootstrap samples:
For each bootstrap replicate $b=1,\ldots,B$ (typically B=200B=200500500), generate a synthetic dataset $\{y_{t_{1}}^{b},\ldots,y_{t_{n}}^{b}\}$ by sampling from the fitted model:
$y_{t_{j}}^{b}\sim\mathcal{N}\left(\mu_{j}(\hat{\boldsymbol{\theta}}),\hat{\sigma}^{2}\right),\quad j=1,\ldots,n,$ (13)
where $\hat{\boldsymbol{\theta}}$ and $\hat{\sigma}^{2}$ are the parameter estimates from the original data.
4.2.2.2 Re-estimate parameters:
For each bootstrap sample $\{y_{t_{1}}^{b},\ldots,y_{t_{n}}^{b}\}$, solve the NLS optimization problem (Eq. 11) to obtain bootstrap parameter estimates $\hat{\boldsymbol{\theta}}^{b}$.
4.2.2.3 Construct confidence intervals:
Use the empirical distribution of $\{\hat{\boldsymbol{\theta}}^{b}:b=1,\ldots,B\}$ to construct confidence intervals. For example, the 95% confidence interval for parameter $\theta_{k}$ is given by the 2.5th and 97.5th percentiles of $\{\hat{\theta}_{k}^{b}:b=1,\ldots,B\}$.
For forecasting, we propagate parameter uncertainty by solving the ODE model for each bootstrap parameter estimate $\hat{\boldsymbol{\theta}}^{b}$ to obtain forecast trajectories $\hat{y}^{b}(t_{n+h})$ for $h$-step-ahead predictions. The 95% prediction interval for $y(t_{n+h})$ is constructed from the 2.5th and 97.5th percentiles of $\{\hat{y}^{b}(t_{n+h}):b=1,\ldots,B\}$.
4.2.3 Computational Implementation
Frequentist inference was implemented using the QuantDiffForecast (QDF) MATLAB toolbox Chowell2024QuantDiffForecast, which provides a comprehensive framework for parameter estimation, uncertainty quantification, and forecasting for ODE models. The QDF toolbox supports multiple optimization algorithms, flexible error structures (normal, Poisson, negative binomial), user-defined ODE systems, and automated bootstrap-based uncertainty quantification. The toolbox outputs parameter estimates with confidence intervals, fitted model trajectories, forecast distributions, and performance metrics (MAE, MSE, WIS, 95% PI coverage).
4.3 Performance Metrics
[…]
4.3.1 Mean Absolute Error (MAE)
MAE measures the average absolute deviation between model predictions and observed data:
$\text{MAE}=\frac{1}{N}\sum_{i=1}^{N}\left|f(t_{i},\hat{\boldsymbol{\theta}})-y_{t_{i}}\right|.$ (14)
MAE provides a direct measure of forecast accuracy on the original data scale. Lower MAE values indicate better point forecast performance. MAE is less sensitive to outliers compared to MSE.
4.3.2 Mean Squared Error (MSE)
MSE measures the average squared deviation between model predictions and observed data:
$\text{MSE}=\frac{1}{N}\sum_{i=1}^{N}\left(f(t_{i},\hat{\boldsymbol{\theta}})-y_{t_{i}}\right)^{2}.$ (15)
MSE penalizes larger errors more heavily than MAE due to the squaring operation, making it more sensitive to outliers. Lower MSE values indicate better fit. The square root of MSE (RMSE) is sometimes reported to return the metric to the original data scale.
4.3.3 Coverage of the 95% Prediction Interval
The 95% PI coverage quantifies the proportion of observed data points that fall within the 95% prediction interval, providing a measure of uncertainty calibration:
$\text{95\% PI Coverage}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(L_{t_{i}}<y_{t_{i}}<U_{t_{i}}),$ (16)
where $L_{t_{i}}$ and $U_{t_{i}}$ are the lower and upper bounds of the 95% prediction interval at time $t_{i}$, and $\mathbf{1}(\cdot)$ is an indicator function equal to 1 if the condition is true and 0 otherwise. Ideally, the coverage should be close to 95%. Coverage significantly below 95% indicates overly narrow prediction intervals (underestimation of uncertainty), while coverage above 95% suggests overly wide intervals (overestimation of uncertainty).
4.3.4 Weighted Interval Score (WIS)
WIS is a proper scoring rule that evaluates the quality of the entire predictive distribution by combining sharpness (interval width) and calibration (penalties for observations outside intervals) (Gneiting2007; bracher2021evaluating). The interval score (IS) for a single prediction interval at level $\alpha$ is:
$\text{IS}_{\alpha}(F,y)=(u-l)+\frac{2}{\alpha}(l-y)\mathbf{1}(y<l)+\frac{2}{\alpha}(y-u)\mathbf{1}(y>u),$ (17)
where $l$ and $u$ represent the $\frac{\alpha}{2}$ and $\left(1-\frac{\alpha}{2}\right)$ quantiles of the forecast distribution $F$, respectively. The IS consists of three components:
4.3.4.1 Sharpness
(ul)(u-l), the width of the central $(1-\alpha)\times 100\%$ prediction interval. Narrower intervals receive lower scores, rewarding precise forecasts.
4.3.4.2 Underprediction penalty
$\frac{2}{\alpha}(l-y)\mathbf{1}(y<l)$, penalizes observations falling below the lower bound $l$, with the penalty proportional to the distance (ly)(l-y).
4.3.4.3 Overprediction penalty
$\frac{2}{\alpha}(y-u)\mathbf{1}(y>u)$, penalizes observations exceeding the upper bound $u$, with the penalty proportional to the distance (yu)(y-u).
To comprehensively evaluate the full predictive distribution, we compute the WIS by averaging interval scores over multiple prediction interval levels $(1-\alpha_{1})<(1-\alpha_{2})<\cdots<(1-\alpha_{K})$ along with the predictive median $\tilde{y}$ (which can be viewed as a prediction interval at level $(1-\alpha_{0})\to 0$):
$\text{WIS}_{\alpha_{0:K}}(F,y)=\frac{1}{K+\frac{1}{2}}\left(w_{0}|y-\tilde{y}|+\sum_{k=1}^{K}w_{k}\,\text{IS}_{\alpha_{k}}(F,y)\right),$ (18)
where $w_{k}=\frac{\alpha_{k}}{2}$ for $k=1,\ldots,K$ and $w_{0}=\frac{1}{2}$. WIS quantifies how close the entire predictive distribution is to the observed data in units on the original data scale. Lower WIS values indicate better forecast performance, balancing sharpness and calibration.
4.3.5 Interpretation
Lower values of MAE, MSE, and WIS indicate better model performance. For 95% PI coverage, values close to 95% indicate well-calibrated uncertainty estimates. Together, these metrics provide a comprehensive assessment of both point forecast accuracy (MAE, MSE), distributional forecast quality (WIS), and uncertainty calibration (95% PI coverage). We report these metrics separately for the calibration period (in-sample fit) and forecasting period (out-of-sample prediction) to distinguish between model fitting and predictive performance.
5 Structural Identifiability
Structural identifiability (SI) analysis determines whether model parameters can, in principle, be uniquely recovered from perfect, noise-free observations of the system outputs, given the model structure and observation scheme (raue2009structural; identifiability_review; stigter2023structural). A parameter is structurally identifiable if its value can be uniquely determined from the input-output relationship defined by the model equations and the set of observable variables. Conversely, a parameter is structurally unidentifiable if multiple distinct parameter values produce identical model outputs, making unique parameter recovery impossible regardless of data quality or quantity.
Structural identifiability is a prerequisite for meaningful parameter estimation: if a parameter is structurally unidentifiable, no amount of data or sophisticated inference methods can recover its true value (practical_identifiability; chowell2023structural). SI analysis is particularly important in compartmental models where only a subset of state variables is observed, as partial observability often leads to identifiability deficits (dankwa2022identifiability). Understanding which parameters are identifiable under different observation scenarios guides experimental design, informs prior specification in Bayesian inference, and helps interpret estimation results.
5.1 Methodology
We conducted symbolic structural identifiability analysis using StructuralIdentifiability.jl (stigter2023structural), a Julia package that employs differential algebra methods to determine identifiability of parameters in ODE models. For each model, we analyzed two scenarios: (1) known initial conditions (ICs), where the initial values of all state variables are assumed to be known exactly, and (2) unknown initial conditions, where ICs are treated as additional unknown parameters to be estimated. The distinction between known and unknown ICs is critical, as identifiability results can differ substantially depending on whether ICs must be inferred from data (identifiability_methods).
For the Lotka–Volterra (LV) model, we examined three observation schemes to assess how partial observability affects identifiability:
5.1.0.1 LV-1 (Both prey and predator observed):
Both $x(t)$ (prey) and $y(t)$ (predator) populations are observed.
5.1.0.2 LV-2 (Predator only):
Only $y(t)$ (predator) is observed; prey population $x(t)$ is unobserved.
5.1.0.3 LV-3 (Prey only):
Only $x(t)$ (prey) is observed; predator population $y(t)$ is unobserved.
For the Generalized Logistic Model (GLM), we analyzed the identifiability when observing incident cases (the time derivative of cumulative cases $C(t)$). For the SEIUR model, we considered the scenario where only incident reported cases (the flow $\kappa\rho E(t)$) are observed, representing typical epidemic surveillance data where only a fraction of infections are detected and reported.
5.2 Results
Table 2 summarizes the structural identifiability results for all models under different observation schemes and initial condition assumptions.
Model Observations Identifiable (unknown IC) Unidentifiable (unknown IC) Identifiable (known IC) Unidentifiable (known IC)
LV-1 $x(t),y(t)$ $\alpha,\beta,\delta,\gamma,x(t),y(t)$ $\alpha,\beta,\delta,\gamma,x(0),y(0)$
LV-2 $y(t)$ only $\alpha,\beta,\gamma,y(t)$ $\delta,x(t)$ $\alpha,\beta,\delta,\gamma,x(0),y(0)$
LV-3 $x(t)$ only $\alpha,\delta,\gamma,x(t)$ $\beta,y(t)$ $\alpha,\beta,\delta,\gamma,x(0),y(0)$
GLM C(t),dC/dtC(t),\,dC/dt C,r,K,pC,\,r,\,K,\,p All states and parameters
SEIUR $\kappa\rho E(t)$ $C,\,\kappa,\,\gamma_{1}$ $S,\,E,\,I,\,U,\,R,\,N,\,\rho$ All states and parameters
Table 2: Structural identifiability results for all models under different observation schemes. The table indicates which parameters and state variables are identifiable or unidentifiable when initial conditions (ICs) are known versus unknown. LV-1, LV-2, and LV-3 refer to Lotka–Volterra scenarios with both species observed, predator only, and prey only, respectively.
5.3 Interpretation
5.3.0.1 Lotka–Volterra Model.
When both predator and prey time series are observed (LV-1), all model parameters ($\alpha,\beta,\gamma,\delta$) are structurally identifiable regardless of whether initial conditions are known or unknown. The coupled dynamics of the two species provide sufficient information to uniquely determine all interaction parameters.
With partial observation, identifiability losses emerge when initial conditions are unknown. When only the predator is observed (LV-2), the prey growth-from-consumption parameter $\delta$ and the prey population trajectory $x(t)$ become unidentifiable. The observed predator dynamics constrain only the product $\delta x(t)$ rather than $\delta$ and $x(t)$ separately, creating a structural non-uniqueness. Similarly, when only the prey is observed (LV-3), the predation rate $\beta$ and the predator population $y(t)$ are unidentifiable because the prey dynamics constrain only the product $\beta y(t)$.
Importantly, when initial conditions are known, all parameters become identifiable even under partial observation scenarios (LV-2 and LV-3). Knowledge of the initial prey and predator populations breaks the symmetry and allows unique recovery of all parameters. This highlights the value of accurate initial condition estimates in ecological field studies.
5.3.0.2 Generalized Logistic Model.
For the GLM, observation of cumulative cases $C(t)$ and incident cases $dC/dt$ yields full structural identifiability of all parameters $(r,p,K)$ under both known and unknown initial conditions. The single-equation structure of the GLM, combined with direct observation of both the state variable and its derivative, ensures that all growth parameters are uniquely determined. This robust identifiability makes the GLM a reliable choice for phenomenological epidemic modeling when the full case trajectory is observed.
5.3.0.3 SEIUR Model.
The SEIUR model exhibits substantial identifiability challenges due to its multi-compartment structure and partial observability. Even when all compartments $(S,E,I,U,R,C)$ are theoretically observable, only the cumulative reported cases $C$, the incubation rate $\kappa$, and the recovery rate $\gamma_{1}$ remain structurally identifiable when initial conditions are unknown. The remaining state variables ($S,E,I,U,R$) and critical parameters including the total population size $N$ and the reporting proportion $\rho$ are structurally unidentifiable.
This identifiability deficit arises because the observed incidence data (new reported cases per day) depend on the product $\kappa\rho E(t)$ rather than on $\kappa$, $\rho$, and $E(t)$ individually. Multiple combinations of reporting proportion $\rho$, exposed population $E(t)$, and total population $N$ can produce identical observed case counts. The unobserved unreported infectious compartment $U(t)$ further compounds the identifiability problem, as the split between reported and unreported infections is not directly constrained by the data.
When initial conditions are known, all SEIUR parameters become identifiable. This underscores the importance of accurate estimates of the initial susceptible and exposed populations (e.g., from seroprevalence surveys or contact tracing data) for reliable parameter inference in epidemic models with underreporting.
5.4 Implications for Inference
The structural identifiability analysis provides critical context for interpreting the empirical parameter estimates and forecasting results presented in the Results section. For the LV model, we expect both Bayesian and Frequentist methods to successfully recover all parameters when both species are observed (LV-1), but parameter estimates may be unreliable or exhibit high uncertainty in the partial observation scenarios (LV-2, LV-3) where structural identifiability is compromised.
For the GLM applied to lung injury and mpox data, the robust structural identifiability suggests that both inference methods should yield well-constrained parameter estimates and reliable forecasts, provided the data quality is sufficient.
For the SEIUR model applied to COVID-19 data in Spain, the limited structural identifiability indicates that several model parameters cannot be uniquely determined from case incidence data alone. We expect parameter estimates for $\rho$, $N$, and the latent compartments to exhibit substantial uncertainty and potential non-uniqueness. In this setting, Bayesian inference may have an advantage by incorporating prior information to regularize the otherwise ill-posed inverse problem, while Frequentist methods may struggle without additional constraints or data sources.
These predictions will be assessed empirically in the Results section, where we compare parameter estimation accuracy, uncertainty quantification, and forecast performance across models and observation scenarios.
Summary of Models, Domains, Data Sources, and Observation Scenarios
As summarized in Table 3, the models, application domains, data sources, and observation scenarios used in this study are presented.
Model Domain Data Source Scenarios
Lotka–Volterra (LV) Ecology Hudson–Bay lynx–hare population Full (predator + prey), Prey-only, Predator-only
Generalized Logistic Model (GLM) Clinical Lung injury progression Full data
Generalized Logistic Model (GLM) Epidemiology 2022 U.S. mpox outbreak Full data
SEIUR Epidemiology COVID-19 incidence in Spain First-wave data
Table 3: The models, application domains, data sources, and observation scenarios used in this study. The datasets span ecological, clinical, and epidemiological systems.
6 Results
In this section, we present parameter estimation results, performance metrics, and convergence diagnostics for all models and datasets analyzed in this study. We organize results by model, with each subsection containing parameter estimates, forecasting performance, and uncertainty quantification for both Bayesian (BFF) and Frequentist (QDF) inference methods. The structural identifiability analysis (Section 5) provides theoretical context for interpreting these empirical results.
6.1 Lotka–Volterra Model
We fitted the LV model to Hudson Bay lynx-hare data under three observation scenarios: (1) both prey and predator observed (LV-1), (2) predator only (LV-2), and (3) prey only (LV-3). These scenarios allow us to assess how partial observability affects parameter recovery and forecasting performance in light of the structural identifiability results presented in Section 5. To enhance the analysis, we consider these three distinct scenarios. First, we present the results of fitting the model to both prey and predator data simultaneously. Next, we focus on the scenario where the model is fitted exclusively to predator data. Finally, we provide the results for the case where only the prey data is used to fit the model.
6.1.1 Both Prey and Predator
In this section, we present the results with a focus on fitting the model to both prey and predator data.
Figure (5) shows that the BFF method is more effective in capturing the dynamics of the data. in particular, the prey population is modeled with a better range of uncertainty, covering all observed data points when BFF is applied. Furthermore, the method excels at capturing the peak of the predator population, particularly the first peak in the predator population, achieving a near-perfect fit.
Table 4), presents significantly different parameter estimates, despite using both datasets simultaneously. in particular, the parameter $\beta$ is estimated to be much lower in BFF compared to QDF, with the value approaching zero. Additionally, the parameter $\delta$ is the only one that shows approximately the same estimate in both methods.
Table 5) confirms the claim that BFF performs much better than QDF.
Table 6) presents the parameter $\hat{R}$ for all the parameters of the model, which is consistently 1.01. MCMC sampling was conducted with 2000 iterations, and these values of $\hat{R}$ ensure that the chains have converged successfully.
$\alpha$ $\beta$ $\delta$ $\gamma$
BFF .40 (.36, .44) .02 (.02, .02) .03 (.03, .03) 1.06 (.95, 1.17)
QDF .55 (.51,.60) .30 (.32,.35) .03 (.02,.03) .84 (.76,.92)
Table 4: Parameter estimates (joint predator–prey). $\beta$ differs across methods, while $\delta$ agrees.
The Figure 5 shows fitting the Hudson-Bay dataset to the LV model, assuming both prey and predator data simultaneously. The first row represents the results obtained using the BFF method, while the second row shows the results from the QDF method. The first column represents the predator population, while the second column corresponds to the prey population.
Figure 5: LV fit with joint predator–prey observations. Using both time series captures the coupled oscillations and aligns peaks and troughs in each species, indicating a well-calibrated interaction structure.
MAEMAE MSEMSE WISWIS $95\%PI$
Predator BFF 4.95 35.93 3.07 100.00
QDF 2.87 13.56 1.88 100.00
Prey BFF 2.24 8.04 1.46 100.00
QDF 3.29 21.96 2.27 90.48
Table 5: Forecast metrics (joint predator–prey). QDF lowers errors for predator; BFF attains perfect prey coverage.
$\alpha$ $\beta$ $\delta$ $\gamma$
$\hat{R}$ (BFF) 1.01 1.01 1.01 1.01
Table 6: Convergence diagnostics (BFF). All $\hat{R}\approx 1.01$.
In the next sections, We are comparing the performance of two different fitting approaches: BFF and QDF. Since we apply both methods separately when fitting only one component of the dataset (either predator or prey), we should expect to generate two figures, one for each method’s fit.
6.1.2 Predator data only
In the following, we present the results of fitting the Hudson-Bay dataset to the LV model, considering only predator data.
$\alpha$ $\beta$ $\delta$ $\gamma$
Predator BFF 1.2(.75,1.9) .03(.01,.06) .02(.01,.03) .36(.22,.62)
QDF .97(.64,1) .02(.02,.06) .025(.02,.036) .46(.4,.7)
Table 7: The parameter estimation while fitting the Hudson-Bay dataset to the LV model, assuming predator data only. The first row represents the results obtained using the BFF method, while the second row shows the results from the QDF method.
Figure 6 below depicts fitting the Hudson-Bay dataset to the Lotka-Volterra model, assuming only predator. The first column represents the results obtained using the BFF method, while the second column shows the results from the QDF
Figure 6: Predator-only fit under the LV model. Calibrating to a single species reproduces coarse predator cycles but leaves prey dynamics unconstrained, reflecting weaker identifiability of interaction terms.
Table 8 below presents the performance metrics, showing that QDF outperforms BFF in terms of MAE, MSE, and WIS, indicating lower error and better predictive performance.
MAEMAE MSEMSE WISWIS $95\%PI$
Predator BFF 3.6 18.07 2.22 100.00
QDF 2.8 19.7 2.33 95.24
Table 8: The performance metrics while fitting the Hudson-Bay dataset to the LV model, assuming predator data only.
$\alpha$ $\beta$ $\delta$ $\gamma$
Predator $\hat{R}$ (BFF) 1 1 1 1
Table 9: The convergence analysis while fitting the Hudson-Bay dataset to the LV model, assuming predator data only.
The results indicate that QDF provides a better fit to predator data alone, capturing key trends in the population more effectively than BFF.
6.1.3 Prey data only
In the following, we present the results of fitting the Hudson-Bay dataset to the LV model, considering only the prey population.
$\alpha$ $\beta$ $\delta$ $\gamma$
Prey BFF .34(.28,.52) .01(0,.05) .04(.02,.05) 1.4(.75,1.9)
QDF .42(.34,.55) .013(.005,.035) .03(.02,.04) 1.2(.78,1.5)
Table 10: The parameter estimation while fitting the Hudson-Bay dataset to the Lotka-Volterra model, assuming prey data only . The first row represents the results obtained using the BFF method, while the second row shows the results from the QDF method.
Figure 7 below illustrate fitting the Hudson-Bay dataset to the Lotka-Volterra model, assuming prey data only. The first column represents the results obtained using the BFF method for prey data only, while the second column shows the results from the QDF method for prey data only.
Figure 7: Prey-only fit under the LV model. Fitting to prey alone matches overall prey amplitude but yields ambiguous predator dynamics, again indicating partial identifiability with single-species data.
MAEMAE MSEMSE WISWIS $95\%PI$
Prey BFF 4.6 27.34 2.8 100.00
QDF 3.3 15.83 2.1 90.48
Table 11: The performance metrics while fitting the Hudson-Bay dataset to the LV model, assuming prey data only.
Table 12 displays the Gelman-Rubin diagnostic ($\hat{R}$) values for BFF, all of which are 1.01, indicating successful MCMC convergence.
$\alpha$ $\beta$ $\delta$ $\gamma$
Prey $\hat{R}$ (BFF) 1.01 1.01 1.01 1.01
Table 12: The convergence analysis while fitting the Hudson-Bay dataset to the LV model, assuming prey data only.
These results suggest that QDF provides a more robust and accurate model for prey population dynamics, outperforming BFF in terms of both error metrics and parameter estimates.
6.1.4 Comparative analysis of parameter estimates ($\alpha,\beta,\delta,\gamma$) for the LV predator-prey model across observation scenarios (prey-only, predator-only, and predator-prey)
Figure 8 below presents a comparative analysis of the parameter estimates ($\alpha,\beta,\delta,\gamma$) for the (LV) predator-prey model. The estimates are shown in three observation scenarios: prey-only, predator-only, and predator-prey. Each point represents the mean estimate, and the horizontal bars denote the associated uncertainty intervals. Logarithmic scaling is applied to better visualize variations across small-magnitude parameter ranges. This comparison highlights how the availability of data influences the identifiability of the parameter and the consistency between the forecasting methodologies.
Figure 8: Parameter comparison across observation schemes. Joint data produce tighter intervals; single-series widens uncertainty.
6.1.5 Performance metrics comparison of MAE, MSE, WIS, and 95% PI for predator-only vs. predator–prey data in the LV model
Figure 9 below evaluates the predictive performance of the BFF and QDF frameworks under two data availability scenarios: (1) when only predator data are observed and (2) when predator and prey data are available. Forecast accuracy is assessed using four metrics: Mean Absolute Error (MAE), Mean Squared Error (MSE), Weighted Interval Score (WIS) and 95% PI. Across all metrics, using full predator-prey data improves model performance compared to relying on predator-only data, underscoring the value of joint observations in dynamic ecological systems.
Figure 9: Predictive accuracy vs. data availability in LV. Across MAE, MSE, WIS, and 95% PI, joint predator–prey observation improves performance for both frameworks compared with predator-only data.
6.1.6 Performance metrics comparison of MAE, MSE, WIS, and 95% PI for prey-only vs. predator–prey data in the LV model
Figure 10 below evaluates the predictive performance of the BFF and QDF frameworks under two data availability scenarios: (1) when only prey data are observed and (2) when both predator and prey data are available. Forecast accuracy is assessed using four metrics: Mean Absolute Error (MAE), Mean Squared Error (MSE), Weighted Interval Score (WIS) and 95% PI. Across all metrics, using full predator-prey data improves model performance compared to relying on prey-only data, underscoring the value of joint observations in dynamic ecological systems.
Figure 10: Effect of adding predator–prey data when starting from prey-only series. Using both species lowers MAE/MSE, reduces WIS, and raises 95% PI for BFF and QDF relative to prey-only fits.
6.1.7 Posterior Histograms of LV Model Parameters $(\alpha,\beta,\delta,\gamma)$ Across Data Scenarios
The posterior histograms of the LV parameters $(\alpha,\beta,\delta,\gamma)$ are shown in Figures (S1)–(S6). These plots compare predator–prey, predator-only and prey-only scenarios under both BFF and QDF. Each histogram summarizes parameter uncertainty with mean, median and 95% confidence interval (CI), providing insight into identifiability and precision across data scenarios.
Specifically, Figures (S1)–(S3) show the BFF results, while Figures (S4)–(S6) present the QDF estimates. Together, they highlight method-specific differences while confirming the model’s ability to recover biologically meaningful values. Narrow, sharply peaked histograms (e.g., $\alpha\approx 0.4\pm 0.04$ in (Figure S1) indicate strong identifiability and low posterior uncertainty. Broad or flat histograms (e.g., predator-only $\beta$ or $\gamma$ in (Figure S2) reflect weak identifiability, these parameters can vary widely while still fitting the data. Comparing BFF vs QDF panels shows that both recover biologically reasonable values, but QDF tends to yield slightly tighter intervals when data coverage is richer.
6.1.8 Comparison of the parameters of the LV model (BFF) using a Normal distribution as the error structure.
Parameter Prior Distribution Lower Bound (LB) Upper Bound (UB)
$\alpha$ normal(0.5, 0.5) T[0,] 0 NA
$\beta$ normal(0.5, 0.5) T[0,] 0 NA
$\delta$ normal(0.5, 0.5) T[0,] 0 NA
$\gamma$ normal(0.5, 0.5) T[0,] 0 NA
Table 13: Prior distributions and bounds for the parameters in the LV model. The priors are truncated normal distributions with mean 0.5 and standard deviation 0.5, constrained to be non-negative (T[0,]).
6.1.9 Comparison of the parameters of the LV model (QDF) using a Normal distribution as the error structure.
Parameter Lower Bound (LB) Upper Bound (UB)
$\alpha$ 0.001 1.0
$\beta$ 0.001 0.5
$\delta$ 0.001 0.5
$\gamma$ 0.001 1.5
Table 14: Parameter ranges for the LV model of (QDF).
6.2 GLM model for Lung Injury
This section summarizes the results of fitting the generalized logistic model (GLM) to the lung-injury time series. We compare model fits, parameter estimates, predictive accuracy (MAE, MSE, WIS), and the coverage of 95% PI. Figure 11 shows that both methods capture the rise and fall of the affected population and align well at peaks and during the decline phase. Intervals are tight around turning points, with small differences between methods.
Figure 11: GLM fits to the lung-injury data using frequentist inference (QDF) and Bayesian inference (BFF); both capture the overall rise–fall pattern.
Table 15 reports $r,p,$ and $k$ estimates; both methods agree closely on central values. The first row represents the results obtained using the BFF method, while the second row shows the results from the QDF method.
r p k
BFF 0.33 (0.3,0.36) 0.98 (0.96,1) 2734.03 (2732.03,2735.82)
QDF 0.32 (0.30,0.35) 1 (0.99,1.02) 2549.5 (2438.9,2662.9)
Table 15: Parameter estimates for the GLM (lung injury): both frameworks produce similar central estimates for $r$, $p$, and $k$.
Table 16 compares MAE, MSE, WIS, and 95% PI, where the frequentist QDF fit achieve lower errors with comparable coverage.
[…]
BFF 11.85 412.29.43 7.15 97.83
QDF 9.42 152.80 5.75 100
Table 16: Performance metrics for the GLM (lung injury): QDF yields lower MAE, MSE, and WIS with comparable 95% PI.
Table 17 summarizes the convergence of the Bayesian fit (all $\hat{R}=1$), indicating good mixing.
r p k
$\hat{R}$ (BFF) 1 1 1
Table 17: Convergence (Bayesian GLM, lung injury): all parameters have $\hat{R}=1$, indicating good MCMC convergence.
6.2.1 Comparison of Parameter Estimates for the GLM Model using Lung Injury data
Figure 12 shows the comparison of $r,p$ and $k$ between the methods, showing similar point estimates and slightly different uncertainty widths. Each point represents the posterior mean, with horizontal lines indicating the 95% PI. The comparison highlights the consistency in parameter estimation across methods while also revealing subtle differences in uncertainty ranges.
Figure 12: Parameter comparison for the GLM (lung injury): $r$, $p$, and $k$ agree across methods, with small differences in interval widths.
6.2.2 Performance metrics comparison of MAE, MSE, WIS, and 95% PI for the GLM Model using lung injury data
Figure 13 visualizes MAE, MSE, WIS, and PI . The results show that QDF consistently achieves lower error rates while maintaining comparable PI coverage, indicating more accurate and reliable forecasts.
Figure 13: Performance comparison for the GLM (lung injury): QDF attains lower errors while preserving PI.
6.2.3 Posterior Histograms of GLM Parameters ($r,k,p$) Across Data Scenarios
The posterior (or sampling) histograms of the GLM parameters $(r,k,p)$ are shown in Figures (S7)–(S8). Each set summarizes the uncertainty of parameter under BFF and QDF, reporting the mean, median, and 95% confidence interval (CI) for each parameter. Together, these panels allow for side-by-side comparison of how the two estimation frameworks explore $(r,k,p)$ across the data scenarios. Both BFF and QDF yield concentrated posteriors, for example, $p\approx 1.0\pm 0.02$ showing that these parameters are well identified.
6.2.4 Comparison of GLM Parameter Settings (BFF) with Normal Error
Parameter Lower Bound (LB) Upper Bound (UB)
$r$ 0.0 NA
pp 0.0 NA
$k$ 0.0 NA
Table 18: Parameter ranges for the GLM model.
6.2.5 Comparison of GLM Parameter Settings (QDF) with Normal Error
Parameter Lower Bound (LB) Upper Bound (UB)
$r$ 0.01 5.0
pp 0.0 2.0
$k$ 0.01 1000000.0
Table 19: Parameter ranges for the GLM model.
6.3 GLM model for mpox epidemic in the USA
In this section, we present the result of fitting the generalized logistic model (GLM) to the 2022 U.S. mpox epidemic.
Figure 14 shows GLM fits to U.S. mpox incidence; both frequentist (QDF) and bayesian (BFF) capture the rapid rise and decline. The first column represents the results obtained using the QDF, while the second column shows the results from the BFF method.
Figure 14: GLM fits to U.S. mpox incidence using frequentist inference (QDF) and bayesian inference (BFF); both reproduce the observed rise–fall pattern.
Table 20 reports parameter estimates $(r,p,k)$ for the mpox GLM; the methods give similar central values with modest interval differences.
r p k
BFF 1.95 (1.76,2.15) 0.84 (0.82,0.85) 29041.01 (27784.85,30164.68)
QDF 2.25 (2.01,2.7) 0.81 (0.80,0.83) 30363.6 (29168.42,31838.14)
Table 20: Parameter estimates for the GLM (mpox): both frameworks yield similar point estimates for $r$, $p$, and $k$ with modest differences in uncertainty.
Table 21 compares MAE, MSE, WIS, and 95% PI; QDF attains lower error metrics with comparable coverage.
[…]
BFF 11.739 399.43 7.175 97.8
QDF 9.45 150.47 5.8 100
Table 21: Performance metrics for the GLM (mpox): QDF achieves lower MAE, MSE, and WIS with coverage near 95% PI, comparable to BFF.
Table 22 summarizes the convergence for the Bayesian fit, with all $\hat{R}=1$ indicating good mixing.
r p k
$\hat{R}$ (BFF) 1 1 1
Table 22: Convergence diagnostics for the Bayesian GLM (mpox): all parameters have $\hat{R}=1$, indicating good MCMC convergence.
6.3.1 Comparison of the parameters estimation of GLM model using Mpox data
Figure 15 compares $(r,p,k)$ between methods, showing close agreement and slightly different uncertainty widths. Each point represents the posterior mean, with horizontal lines indicating the 95% PI. The comparison highlights the consistency in parameter estimation across methods while also revealing subtle differences in uncertainty ranges.
Figure 15: Parameter comparison for the GLM (mpox): $r$, $p$, and $k$ estimates are consistent across QDF and BFF, with slightly different interval widths.
6.3.2 Performance metrics comparison of MAE, MSE, WIS, and 95% PI for the GLM Model using mpox data
Figure 16 visualizes the error and coverage metrics. Both methods produced similar results for all four evaluation metrics: MAE, MSE, WIS, and 95% PI. The MAE values are nearly identical, indicating comparable point forecast accuracy. MSE is slightly lower for QDF, suggesting marginally better handling of larger errors, while The (WIS) values are very close, reflecting similar quality in predictive intervals. Both approaches achieved nearly identical 95% PI, indicating that their uncertainty quantification is equally well-calibrated. Overall, both methods performed comparably for the mpox data in the GLM model.
Figure 16: Performance comparison for the GLM (mpox): low errors and near-nominal 95% PI indicate good calibration on this dataset.
6.3.3 Posterior Histograms of GLM Parameters ($r,k,p$) for Mpox Data
The posterior (BFF) and sampling (QDF) histograms of the GLM parameters $(r,k,p)$ for the mpox incidence data are shown in Figures (S9)–(S10). Each panel summarizes the uncertainty with the mean, median, and 95% CI, allowing side-by-side comparison of BFF and QDF on the same dataset. Figures(S10) You will see $r$ concentrated roughly between $\sim 2$ and $\sim 3.3$, $p$ around $\sim 0.80$$0.85$, and $k$ centered near $\sim 3.0\times 10^{4}$ (tight unimodal shapes $\rightarrow$ with tight identifiability). Figures(S9) The printed stats on the panels show very tight posteriors (e.g., $p\approx 0.98,k\approx 2734,r\approx 0.33$). The GLM parameters are clearly identified under both methods; verify the Figure (S9) labeling in your source before submission.
6.3.4 GLM Parameter Settings for Mpox Data (BFF, Normal Error)
Parameter Lower Bound (LB) Upper Bound (UB)
$r$ 0.0 NA
pp 0.0 NA
$k$ 0.0 30000
Table 23: Parameter ranges for the GLM on mpox data under BFF with Normal error.
6.3.5 GLM Parameter Settings for Mpox Data (QDF, Normal Error)
Parameter Lower Bound (LB) Upper Bound (UB)
$r$ 0.01 5.0
pp 0.0 2.0
$k$ 0.01 1000000.0
Table 24: Parameter ranges for the GLM on mpox data under QDF with Normal error.
6.4 SEIUR model of COVID-19 cases in Spain during the first wave
In this section, we present the parameter estimation, performance metrics, and convergence analysis obtained by fitting the COVID-19 first-wave data from Spain to a compartmental SEIUR (Susceptible-Exposed-Infectious-Recovered model).
Figure 17 shows the SEIUR fits of the first wave in Spain; both frameworks track the rise and decline in incidence.
Figure 17: SEIUR fit to Spain’s first COVID-19 wave using QDF and BFF; both reproduce the observed rise–fall in incidence.
Table 25 reports $(\beta_{0},\beta_{1},q_{1},\rho,\kappa,\gamma_{1})$; estimates are comparable between methods with some differences in $\rho$ and $q_{1}$.
$\beta_{0}$ $\beta_{1}$ $q_{1}$ $\rho$ $\kappa$ $\gamma_{1}$
BFF 2.18 (1.6, 3.09) 1.69 (1.14, 2.24) 0.88 (0.64, 1.41) 0.6 (0.28, 0.92) 2.39 (1.23, 3.73) 1.78 (1.22, 2.59)
QDF 1.93 (1.18, 1.96) 1.35 (0.54, 1.38) 1.99 (1.13, 2.00) 0.96 (0.72, 1.00) 1.62 (0.61, 1.75) 1.44 (0.65, 1.47)
Table 25: SEIUR parameter estimates $(\beta_{0},\beta_{1},q_{1},\rho,\kappa,\gamma_{1})$ for Spain first wave under BFF and QDF.
Table 26 compares MAE, MSE, WIS, and 95% PI; BFF attains lower errors and similar coverage than QDF.
[…]
BFF 294.57 210754.1222 211.303 94
QDF 352.2978 311053.5723 250.4568 92
Table 26: SEIUR performance metrics (MAE, MSE, WIS, 95% PI): BFF achieves lower errors with similar coverage relative to QDF.
Table 27 lists convergence for the Bayesian fit; $\hat{R}$ values indicate acceptable mixing overall.
$\beta_{0}$ $\beta_{1}$ $q_{1}$ $\rho$ $\kappa$ $\gamma_{1}$
$\hat{R}$ (BFF) 1.18 1.22 1.17 1 1.07 1.21
Table 27: Convergence diagnostics for the bayesian SEIUR fit; $\hat{R}$ values indicate acceptable chain mixing.
6.4.1 Comparison of the parameters estimation of SEIUR model
Figure 18 compares the parameters ($\beta_{0}$, $\beta_{1}$, $q_{1}$, $\rho$, $\kappa$ and $\gamma_{1}$ )estimates between methods, highlighting a larger uncertainty for several rates under BFF.
Figure 18: Parameter comparison for SEIUR (Spain): cross-method estimates for $(\beta_{0},\beta_{1},q_{1},\rho,\kappa,\gamma_{1})$ with differing uncertainty widths.
6.4.2 Performance metrics comparison of MAE, MSE, WIS, and 95% PI for the SEIUR model
Figure 19 visualizes MAE, MSE, WIS, and PI; BFF is best in all error metrics with similar coverage.
Figure 19: SEIUR performance comparison (Spain): BFF shows lower MAE/MSE/WIS with similar 95% PI.
6.4.3 Posterior Histograms of SEIUR Parameters ($\beta_{0},\beta_{1},q_{1},\rho,\kappa,\gamma_{1},N$)
Figures S11S12 show the distributions of the SEIR parameters under BFF and QDF. Each panel summarizes the uncertainty with the mean, median, and 95% CI) facilitating the method-by-method comparison. Figure S11 (BFF): Posteriors are moderately wide for several rates, indicating only partial identifiability from the data alone. Examples printed on the graphs: $\beta_{0}$ median $2.18$ with 95% CI $[1.60,3.09]$; $\beta_{1}$ median $1.69$ $[1.14,2.48]$; $q$ median $0.88$ $[0.64,1.41]$; $\rho$ median $0.60$ $[0.28,0.92]$; $\kappa$ median $2.39$ $[1.23,3.73]$; $\gamma$ median $1.78$ $[1.22,2.59]$. These wide intervals highlight the coupling among parameters and the limited information in the observed outputs. Figure S12 (QDF): Distributions appear more concentrated for several parameters (narrower support on the x-axis, unimodal), consistent with tighter point estimates under QDF. Use it in conjunction with Figure S11 to contrast where the BFF uncertainty bands are widest. BFF histograms are relatively broad for multiple rates, reflecting partial identifiability, while QDF histograms show sharper concentrations, indicating more stable estimates. This contrast underscores where additional data or reparameterization could improve identifiability.
6.4.4 Comparison of the parameters of the SEIR model (BFF) using a Normal error structure
Parameter Lower Bound (LB) Upper Bound (UB)
$\beta_{0}$ 0.0 NA
$\beta_{1}$ 0.0 NA
$q1$ 0.0 NA
$\rho$ 0.0 NA
$\kappa$ 0.0 NA
$\gamma_{1}$ 0.0 NA
NN 0.0 47000000
Table 28: Parameter ranges for the SEIUR model under BFF with Normal error.
6.4.5 Comparison of the parameters of the SEIUR model (QDF) using a Normal error structure
Parameter Lower Bound (LB) Upper Bound (UB)
$\beta_{0}$ 0.001 2.0
$\beta_{1}$ 0.001 2.0
$q1$ 0.001 2.0
$\rho$ 0.001 1.0
$\kappa$ 0.001 2.0
$\gamma_{1}$ 0.001 4.0
NN 50000 1000000000.0
Table 29: Parameter ranges for the SEIUR model under QDF with Normal error.
7 Discussion
This comparative analysis of Bayesian and Frequentist inference across ecological, clinical, and epidemiological systems reveals a clear practical distinction: Bayesian methods (BFF) deliver superior uncertainty quantification and diagnostic transparency, while Frequentist methods (QDF) often achieve lower point-error metrics with greater computational efficiency. The relative performance of each framework depends critically on the richness of the data, the complexity of the model, and the degree of structural identifiability. Our findings demonstrate that neither approach universally dominates; instead, the optimal choice depends on the specific characteristics of the modeling problem.
7.1 Model-Specific Performance
7.1.0.1 Lotka–Volterra Model.
For the LV model, the performance varied substantially in all three observation scenarios. When both prey and predator populations were observed (LV-1), both methods successfully recovered parameter values, with QDF achieving slightly lower point errors for predator dynamics (MAE 2.87 vs. 4.95) (Table 5) but BFF achieved perfect prediction interval coverage for prey (100% vs. 90.48%) (Table 5). Under partial observation with prey-only data (LV-3), QDF produced a sharper prediction of points , reducing MAE from 4.60 (BFF) to 3.30, MSE from 27.34 to 15.83 and WIS from 2.80 to 2.10 (Table 11). However, BFF maintained a superior calibration with 100% PI compared to QDF 90.48%. Convergence diagnostics confirmed stable MCMC sampling for BFF, with all LV parameters achieving $\hat{R}\approx 1.01$ (Table 6), indicating reliable posterior inference despite partial observability.
The structural identifiability analysis (Section 5) explains these patterns: full observation of both species enables unique parameter recovery, while partial observation creates identifiability deficits that neither inference framework can overcome without known initial conditions. The empirical results align with the SI predictions, demonstrating that observability fundamentally constrains inference quality regardless of the statistical method employed.
7.1.0.2 Generalized Logistic Model (Lung Injury and Mpox).
For GLM applied to lung injury and mpox outbreaks, QDF consistently outperformed BFF in point forecast accuracy while maintaining comparable or superior uncertainty calibration. In the mpox epidemic, QDF achieved lower MAE (9.45 vs. 11.74), MSE (150.47 vs. 399.43), and WIS (5.8 vs. 7.18), with perfect PI coverage (100% vs. 97.8%) (Table 21). BFF demonstrated excellent convergence ($\hat{R}=1$) (Table 17) and (Table 22) for all parameters $(r,p,K)$, confirming that the posterior was well-sampled despite the less favorable performance metrics. The lung injury results exhibited the same qualitative pattern, with QDF achieving lower point errors while maintaining well-calibrated prediction intervals.
The robust structural identifiability of the GLM (Section 5) explains why both methods succeeded in parameter recovery. The single-equation structure combined with direct observation of incident cases ensures that all parameters $(r,p,K)$ are uniquely identifiable under both known and unknown initial conditions. In such well-identified settings, the computational efficiency and optimization focus of QDF provide advantages for point estimation, while BFF’s probabilistic framework incurs additional computational cost without substantial gains in forecast accuracy.
7.1.0.3 SEIUR Model (COVID-19 in Spain).
In contrast to the GLM results, BFF outperformed QDF in all metrics for the SEIUR model: MAE 294.57 vs. 352.30, MSE 210,754 vs. 311,054, WIS 211.30 vs. 250.46, and PI coverage 94% vs. 92% (Table 26). This reversal of relative performance reflects the increased complexity of the model, latent compartments, and limited structural identifiability of the SEIUR framework. SI analysis (Section 5) demonstrates that when only incident reported cases are observed, most parameters and state variables ($S$, $E$, II, $U$, $R$, NN, $\rho$) are structurally unidentifiable when initial conditions are unknown. Only cumulative cases $C$ incubation rate $\kappa$, and recovery rate $\gamma_{1}$ remain identifiable.
In this challenging setting, Bayesian inference benefits from prior regularization that constrains the otherwise ill-posed inverse problem. The posterior distribution incorporates both the likelihood of the data and prior information, enabling more stable parameter estimation and forecast generation despite fundamental identifiability limitations. Frequentist optimization, on the contrary, struggles with the high-dimensional, poorly constrained parameter space, leading to less reliable point estimates and wider prediction intervals.
7.2 Role of Structural Identifiability
The structural identifiability analysis provides critical insight into why inference performance varies across models and observation scenarios. For the LV model, global identifiability requires observation of both prey and predator; partial observation creates parameter non-uniqueness that neither BFF nor QDF can resolve (Table 2). For the GLM, robust identifiability under all observation schemes ensures reliable inference for both frameworks. For the SEIUR model, severe identifiability deficits constrain parameter recovery regardless of the inference method, though Bayesian priors provide some regularization benefit.
These findings underscore that structural identifiability sets fundamental limits on what can be learned from the data. No statistical framework, no matter how sophisticated, can uniquely recover structurally unidentifiable parameters. SI analysis should therefore precede empirical modeling to clarify what questions the data can and cannot answer.
7.3 Practical Implications
Our results provide practical guidance for selecting between Bayesian and Frequentist inference frameworks:
7.3.0.1 Use Frequentist methods (QDF)
when: (1) data are abundant and of high quality, (2) models are structurally identifiable, (3) computational speed is critical, (4) the primary objective is minimizing point forecast error, and (5) model complexity is low to moderate. The GLM applications exemplify these conditions.
7.3.0.2 Use Bayesian methods (BFF)
when: (1) data are sparse, noisy, or partially observed, (2) models contain latent states or unobserved compartments, (3) structural identifiability is limited, (4) comprehensive uncertainty quantification is essential, (5) prior information is available to constrain inference, and (6) diagnostic assurance via convergence metrics is important. The SEIUR application exemplifies these conditions.
7.3.0.3 Improve both methods
by: (1) increasing observational coverage (e.g., measuring both prey and predator rather than one species), (2) obtaining accurate estimates of initial conditions, (3) incorporating auxiliary data sources to break parameter non-uniqueness, and (4) using SI-guided experimental design to ensure parameters of interest are identifiable.
When practical identifiability is limited or interval calibration is critical, BFF is advantageous because posterior sampling naturally captures parameter and predictive uncertainty while providing convergence assurance (e.g., GLM $\hat{R}=1$; LV $\hat{R}\approx 1.01$). When the priority is minimizing point forecast error and data are informative, QDF can be preferable (e.g., LV prey-only; GLM mpox), though prediction intervals may be under-calibrated relative to BFF (e.g., LV prey-only coverage 90.48% vs. 100%). Across all systems, adding informative observations improves both methods (LV predator-prey vs. single-series), and SI analysis clarifies when further accuracy gains are impossible without additional data or constraints.
7.4 Limitations and Future Directions
This study standardized model structures, observation schemes, and error distributions to isolate the effects of the inference framework. Although this design enables for clean comparison, several limitations warrant consideration. First, we assumed normal error structures for both methods; alternative distributions (e.g., negative binomial for count data) might shift relative performance. Second, we analyzed each model independently; hierarchical or multi-level structures could favor Bayesian inference more strongly. Third, our SI analysis assumed perfect, noise-free data; practical identifiability under realistic noise levels merits further investigation.
Future research could explore hybrid inference strategies that combine the computational efficiency of optimization-based methods with the uncertainty quantification strengths of Bayesian sampling. The SI-guided experimental design could determine the optimal measurement strategies before data collection. Incorporating machine learning approaches such as neural ODEs giampiccolo2024hnode or physics-informed neural networks may offer complementary advantages. Extending the comparison to spatially explicit models, agent-based simulations, and stochastic epidemic frameworks would broaden the applicability of these findings.
8 Conclusion
Mathematical models based on ordinary differential equations are essential tools for understanding and predicting dynamics in ecological and epidemiological systems. The reliability of model-based inferences is critical to accurate parameter estimation and proper assessment of parameter identifiability. This study provides a systematic comparison of Bayesian and Frequentist inference frameworks across three biological models LV predator-prey dynamics, the generalized logistic model, and the SEIUR epidemic model applied to four real world datasets spanning ecology and epidemiology.
Our analysis demonstrates that neither Bayesian nor Frequentist methods universally outperform the other; instead, the optimal framework depends on data characteristics, model structure, and the degree of structural identifiability. Frequentist methods (implemented via QDF) excel when data are abundant, models are well identified, and computational efficiency is prioritized, as demonstrated by superior point forecast accuracy for the GLM applications to lung injury and mpox outbreaks. Bayesian methods (implemented via BFF) excel when data are sparse or noisy, models contain latent compartments, and comprehensive uncertainty quantification is essential, as demonstrated by superior performance for the SEIUR model of COVID-19 in Spain.
Structural identifiability analysis emerged as a critical prerequisite for meaningful inference. Our symbolic identifiability results using StructuralIdentifiability.jl revealed that observability fundamentally constrains parameter recovery: full observation of both predator and prey populations in the LV model enabled complete parameter identification, while partial observation created identifiability deficits that neither statistical framework could overcome without known initial conditions. Similarly, the SEIUR model’s limited identifiability when only reported cases are observed explains why both methods struggled with parameter uncertainty, although Bayesian priors provided some regularization benefit.
Integration of structural identifiability analysis with empirical inference provides a principled workflow for biological modeling. First, conduct a SI analysis to determine which parameters can theoretically be recovered from the planned observations. Second, select an inference framework matched to data characteristics: use Frequentist methods when identifiability is strong and data are rich; use Bayesian methods when identifiability is weak or data are limited. Third, interpret the empirical results in light of SI constraints, recognizing that no statistical method can uniquely recover structurally unidentifiable parameters.
For practitioners selecting between inference frameworks, we recommend: (1) Use Frequentist methods when minimizing point prediction error and computational cost is paramount, data quality is high, and models are structurally identifiable. (2) Use Bayesian methods when uncertainty quantification is essential, prior information is available, models contain unobserved states, or identifiability is limited. (3) Improve both methods by improving observational coverage, obtaining accurate initial condition estimates, and using SI-guided experimental design.
This comparative study advances our understanding of how statistical inference frameworks perform in biological modeling contexts and provides practical guidance for method selection. By aligning the inference approach with problem characteristics and grounding interpretation in identifiability theory, modelers can achieve reliable, reproducible, and scientifically defensible results that inform ecological management and public health decision-making.
Appendix A Supplementary Material
A.1 LV-Model (BFF) Histograms
Figure S1: Histograms for predator and prey data of $(\alpha,\beta,\delta,\gamma)$ Parameters
Figure S2: Histograms for predator data only of $(\alpha,\beta,\delta,\gamma)$ Parameters
Figure S3: Histograms for prey data only of $(\alpha,\beta,\delta,\gamma)$ Parameters
A.2 LV-model (QDF) Histograms
Figure S4: Histograms for predator-prey data of $(\alpha,\beta,\delta,\gamma)$ Parameters
Figure S5: Histograms for predator data only of $(\alpha,\beta,\delta,\gamma)$ Parameters
[…]
A.3 GLM-Model-Lung Injury BayesianFitForecast (BFF) Histograms
Figure S7: Histograms for GLM data of r, k, and p Parameters
A.4 GLM-Model-Lung Injury (QDF) Histograms
Figure S8: Histograms r,p, and k Parameters
A.5 GLM-Model-Mpox (BFF) Histograms
Figure S9: Histograms for GLM data of r, k, and p Parameters
A.6 GLM-Model-Mpox (QDF) Histograms
Figure S10: Histograms for GLM data of r, k, and p Parameters
A.7 SEIUR-Model (BFF) Histograms
Figure S11: Histograms ($\beta_{0}$, $\beta_{1}$, $q_{1}$, $\rho$, $\kappa$, and $\gamma_{1}$ ) Parameters
A.8 SEIUR-Model (QDF) Histograms
Figure S12: Histograms for ($\beta_{0}$, $\beta_{1}$, $q_{1}$, $\rho$, $\kappa$, and $\gamma_{1}$ ) Parameters