### v2

# Comparing Bayesian and Frequentist Inference in Biological Models: A Comparative Analysis of Accuracy, Uncertainty, and Identifiability

Abstract
Mathematical models are widely used to study ecological and epidemiological systems, but the accuracy of resulting inferences and forecasts depends on the estimation framework. This study compares Bayesian and Frequentist approaches across three biological models using four datasets: the Lotka–Volterra (LV) predator–prey system, a generalized logistic model (GLM) applied to lung injury and mpox data, and SEIUR epidemic model for COVID-19. To ensure a fair comparison, both approaches were implemented with a normal error structure. We first examined structural identifiability to determine which parameters can be recovered in principle. We then evaluated practical identifiability and fitting performance using four metrics: mean absolute error (MAE), mean squared error (MSE), prediction interval (PI) coverage, and weighted interval score (WIS). For the Lotka–Volterra model, we studied three types of observations: prey only, predator only, and both species together. The Frequentist workflow, implemented through QuantDiffForecast, uses nonlinear least squares and parametric bootstrap to quantify uncertainty. The Bayesian workflow, implemented through BayesianFitForecast, uses Hamiltonian Monte Carlo sampling to obtain posterior distributions and diagnostic measures. Our results show that Frequentist inference performs well when the system is fully observed data, as in the GLM or in the Lotka–Volterra model when both species are observed. Bayesian inference performs better when uncertainty is high and observations are limited or indirect, as seen in the SEIUR epidemic model. The identifiability analysis helps explain these differences by showing how observability shapes the reliability of parameter recovery and uncertainty quantification.
1 Introduction
Mathematical models based on ordinary differential equations (ODEs) are essential tools for understanding and forecasting dynamics in ecology and epidemiology [1, 2, 3]. Model-based forecasts have been instrumental in managing major public health crises over the past decade. During the COVID-19 pandemic, forecasts guided resource allocation and social distancing policies [4, 5, 6, 7, 8, 9, 91].
The U.S. CDC’s FluSight Challenge used models to optimize influenza vaccine distribution and public health messaging [10, 11, 12]. During the West African and Democratic Republic of the Congo (DRC) Ebola outbreaks, models predicted transmission patterns and evaluated intervention effectiveness [13, 14, 15, 16, 17, 18]. More recently, forecasting models for mpox predicted its spread and assessed containment measures [19, 20, 21, 22].
In ecology, predator–prey models help predict population dynamics and inform conservation strategies [23]. These applications demonstrate that reliable forecasts depend on accurate parameter estimation [24, 25, 26]. However, the ability to recover meaningful parameters from data critically depends on identifiability, that is, whether model parameters can be uniquely determined from available observations [27, 28, 29].
Identifiability can fail when data are sparse, models are over-parameterized, or parameters are strongly correlated, leading to misleading inferences and unreliable forecasts. Recent work has shown that factors such as vaccination behavior, immunity duration, and data completeness shape both epidemic trajectories and the reliability of parameter estimates [30]. Understanding how identifiability interacts with estimation frameworks is essential to ensure that model-based forecasts are interpretable and reproducible. Structural or practical identifiability limitations can cause two inference frameworks to produce different results even under identical models and data.
Two estimation paradigms dominate the field: Bayesian and Frequentist methods. Frequentist methods typically calibrate ODE models by optimizing a likelihood function or minimizing an objective function such as the sum of squared differences between observed and predicted values [31, 32, 33, 34]. These methods use algorithms such as gradient descent or the Levenberg–Marquardt algorithm, assume specific distributions for measurement errors (e.g., Gaussian, Poisson, or Normal), and quantify uncertainty through bootstrapping techniques [35, 36, 37, 38, 39, 40, 41, 42]. Frequentist methods are computationally efficient and often perform well when data are abundant and high quality [43, 25]. The QuantDiffForecast (QDF) toolbox implements this workflow for fitting models and generating predictions with quantified uncertainty [25].
Bayesian methods, in contrast, apply Bayes’ theorem to combine prior distributions of parameters with the likelihood of observed data, producing posterior distributions that explicitly incorporate uncertainty [44, 45, 46, 47, 48, 49, 50, 51]. These methods typically use Markov Chain Monte Carlo (MCMC) algorithms to approximate posterior distributions, providing comprehensive measures of parameter uncertainty and prediction intervals [52, 53, 54, 55, 56, 57]. Bayesian methods can better navigate complex parameter spaces, avoid local optima, and handle incomplete or noisy data [58, 59, 60]. Tools such as Stan facilitate Bayesian estimation and forecasting, allowing rigorous uncertainty quantification and model validation [61, 62, 63, 64, 65].
The BayesianFitForecast (BFF) toolbox implements this workflow with diagnostics such as the Gelman–Rubin $\hat{R}$ statistic [26]. Despite extensive use of both paradigms, prior comparisons often vary models, likelihoods, or preprocessing between methods, making it difficult to attribute observed differences to the estimation framework itself [66]. To address this gap, we conduct a controlled comparison of Bayesian and Frequentist inference under standardized conditions: the same models, the same normal error structure, and harmonized data preprocessing.
We analyze three systems and four datasets representing different levels of complexity and observability: the Lotka–Volterra (LV) predator–prey model (Hudson Bay lynx–hare data), a generalized logistic model (GLM) for lung injury and the 2022 United States mpox outbreak, and an SEIUR epidemic model for the first COVID-19 wave in Spain [67, 68, 69]. For the Lotka–Volterra model, we analyze three observation scenarios (prey only, predator only, and both simultaneously) to assess how partial observability affects parameter recovery.
We integrate structural identifiability analysis to determine which parameters can theoretically be recovered from the data, separating fundamental data limitations from algorithmic limitations [70, 71, 72]. We then evaluate practical identifiability and forecasting performance using four metrics: mean absolute error (MAE), mean squared error (MSE), 95% prediction interval (PI) coverage, and weighted interval score (WIS) [73, 74, 75, 76, 77].
Our objectives are to (i) compare prediction accuracy across Bayesian and Frequentist inference, (ii) assess uncertainty calibration and diagnostics, and (iii) relate observed performance to structural versus practical identifiability under full versus partial observation. This provides practical guidance on when each paradigm is preferable based on data richness, observability, and uncertainty quantification needs [26].
The remainder of this paper is organized as follows. Section 2 describes the four datasets analyzed in this study. Section 3 presents the mathematical models employed. Section 4 details the Bayesian and Frequentist inference methodologies. Section 5 presents the structural identifiability analysis. Section 6 reports the empirical results. The Discussion synthesizes our findings, and the Conclusion summarizes the main contributions and implications for biological modeling practice.
2 Data
We analyzed four datasets spanning ecological and epidemiological systems, each selected to represent different levels of data richness, temporal resolution, and observability (Table 1; Figure 1). All datasets consist of time-series observations at discrete time points $t_{n}$, where $n=1,\ldots,N$ denotes the observation index. The datasets include both population counts (for the ecological system) and incident case counts (for epidemic systems). Data preprocessing was standardized across all datasets to ensure comparability between inference methods; specific preprocessing steps are summarized at the end of this section.
2.1 Hudson Bay Lynx–Hare Data
The Hudson Bay lynx–hare dataset consists of annual counts of Canadian lynx (Lynx canadensis) and snowshoe hare (Lepus americanus) populations from 1900 to 1920, derived from Hudson’s Bay Company pelt return records [76, 77] . These records serve as a proxy for population abundance, with pelt counts reflecting relative population size. The dataset contains $N=21$ yearly observations for both prey (hare) and predator (lynx) populations. This dataset is particularly valuable for assessing parameter identifiability in predator–prey dynamics because it provides simultaneous observations of both interacting species over multiple population cycles. The data have been widely used as a benchmark for testing ecological models and parameter estimation methods. No smoothing, interpolation, or imputation was applied to the original data.
2.2 Lung Injury Data (EVALI)
This dataset comprises weekly incident case counts of electronic-cigarette or vaping product use-associated lung injury (EVALI) reported in the United States during 2019 [78, 79, 80] . The outbreak was first identified in mid-June 2019, with cases rapidly escalating through the summer and fall before declining by November 2019. We analyzed $N\approx 21$ weeks of data spanning mid-June to early November 2019, as reported through the Centers for Disease Control and Prevention (CDC) Morbidity and Mortality Weekly Report (MMWR) surveillance system. Case definitions followed CDC criteria for confirmed and probable EVALI cases. The data were aggregated by epidemiological week (MMWR week) with no backfill corrections applied beyond those included in the official CDC reports. This dataset represents a short-duration outbreak with a clear peak, making it suitable for testing phenomenological growth models under well-defined epidemic dynamics.
2.3 Mpox Data (United States, 2022)
We analyzed weekly incident confirmed or probable mpox cases in the United States during the 2022–2023 outbreak as reported in [81, 82]. The outbreak began in May 2022 and peaked in August 2022 before declining through early 2023. Data were compiled from CDC national surveillance reports aggregated by MMWR week. Case definitions followed CDC guidelines for confirmed (laboratory-confirmed orthopoxvirus with specimen typed as monkeypox virus) and probable cases (epidemiologically linked to a confirmed case without laboratory confirmation). The temporal window analyzed corresponds to the public reporting period documented in CDC surveillance dashboards (see references for access dates). This dataset provides an example of an emerging infectious disease outbreak with intensive surveillance and public health response, allowing evaluation of model performance during rapidly evolving epidemic conditions.
2.4 COVID-19 Data (Spain, First Wave)
This dataset consists of daily incident laboratory-confirmed COVID-19 cases reported during the first epidemic wave in Spain from February through May 2020 [83, 84] . Data were compiled from official reports by the Spanish Ministry of Health (Ministerio de Sanidad), with case dates based on report date rather than symptom onset date due to data availability constraints during the early pandemic response. The first wave in Spain exhibited rapid exponential growth followed by a peak in late March 2020 and subsequent decline following implementation of strict non-pharmaceutical interventions including a national lockdown. We restricted analysis to the first wave period to avoid complications from changing surveillance protocols, intervention policies, and emergence of new variants that occurred in later waves. Known reporting artifacts explicitly flagged in official data releases were excluded. This dataset represents a large-scale epidemic with significant underreporting and latent compartments (exposed and unreported infectious individuals), making it ideal for assessing inference methods under partial observability and model complexity.
Table 1: Summary of datasets analyzed in this study, including data source, temporal coverage, reporting frequency, and number of observations for each system.
Dataset Source & Period Frequency Observations (N)
Hudson Bay Lynx–Hare Hudson’s Bay Company pelt returns (1900–1920) Annual 21
Lung Injury (EVALI, US) CDC MMWR reports (mid-Jun–Nov 2019) Weekly $\approx$21
Mpox Epidemic (US, 2022) CDC surveillance (May 2022–early 2023) Weekly Variable
COVID-19 (Spain, 1st Wave) Spanish Ministry of Health (Feb–May 2020) Daily Variable
2.5 Preprocessing summary:
All datasets were processed using a standardized protocol to ensure comparability across inference methods. The preprocessing steps were as follows:
  1. 1.
    Temporal aggregation: The Hudson Bay lynx–hare data were used as published without aggregation. Lung injury and mpox data were aggregated at the weekly level (MMWR epidemiological week) as reported in official CDC surveillance. COVID-19 data for Spain were analyzed at the daily level as reported by the Spanish Ministry of Health.
  2. 2.
    Data type: The Hudson Bay dataset represents population abundance (pelt counts), while all epidemic datasets (lung injury, mpox, COVID-19) represent incident case counts per reporting period. For the generalized logistic model (GLM), cumulative case counts were derived by summing incident counts, and model predictions were compared to incident data via the time derivative of cumulative cases.
  3. 3.
    Missing values: No missing values were present in the final temporal windows used for model fitting. No imputation was performed.
  4. 4.
    Smoothing: No smoothing was applied to the data used for parameter estimation or forecasting. Figures may display smoothed overlays for visualization purposes only; these smoothed curves were not used in model fitting.
  5. 5.
    Temporal windows: Modeling windows for calibration and forecasting were selected based on epidemic phase and data availability. Specific windows for each dataset are described in the Results section.
  6. 6.
    Data quality control: For COVID-19 data, known reporting artifacts explicitly flagged in official Spanish Ministry of Health releases were excluded. For all other datasets, data were used as published in official surveillance reports without additional corrections.
Figure 1: Time series of ecological and epidemic population dynamics. The Hudson Bay Lynx–Hare dataset shows annual prey and predator abundance (1900–1920), while the remaining panels depict weekly or daily reported cases from major disease outbreaks: Lung Injury (EVALI, US, 2019), Mpox (US, 2022), and COVID-19 (Spain, 2020).
3 Models
In this study, we employed three compartmental biological and epidemiological models to comprehensively assess the parameter identifiability of our two approaches: (i) the Lotka–Volterra model, describing predator-prey interactions and population feedback mechanisms; (ii) the Generalized Logistic Model, capturing flexible epidemic growth through nonlinear case dynamics; and (iii) the SEIUR model, explicitly tracking susceptible, exposed, infectious (reported and unreported), and recovered populations to account for both observed and hidden transmission. These models were selected to represent a progression in model complexity and data observability: from a fully observed ecological system (LV), to a parsimonious single-equation epidemic model (GLM), to a multi-compartment latent-state model with partial observability (SEIUR). This progression enables a systematic comparison of Bayesian and Frequentist inference performance across increasing levels of model structure and latent uncertainty.
3.0.0.1 Lotka-Volterra model
The Lotka–Volterra equations (Lotka 1925; Volterra 1926, 1927) consist of ODEs describing the population dynamics of two interacting species: one predator and one prey. Let $x$ denote the prey population and $y$ the predator population at time $t$. Volterra modeled the temporal dynamics of these populations as follows:
$\frac{dx}{dt}=\alpha x-\beta xy,\quad\frac{dy}{dt}=-\gamma y+\delta xy,$ (1)
with initial conditions
$x(0)=x_{0},\quad y(0)=y_{0},$ (2)
where $\alpha$ is the intrinsic growth rate of the prey population, $\beta$ is the predation rate coefficient, $\gamma$ is the mortality rate of the predator population, and $\delta$ is the predator’s growth rate per prey consumed. The initial conditions $x_{0}$ and $y_{0}$ specify the starting population sizes of prey and predator, respectively. The Lotka–Volterra model serves as a classical benchmark in population ecology, offering a well-understood test case for examining parameter identifiability and uncertainty propagation under full versus partial observability.
xxyyIntrinsic growth $\alpha x$Loss due to predation $\beta xy$Growth from consuming prey $\delta xy$Natural mortality $\gamma y$
Figure 2: Lotka–Volterra predator–prey diagram. circles represent the prey $x$ and predator $y$ populations. The self-loop on $x$ indicates intrinsic growth at rate $\alpha$. Curved arrows between $x$ and $y$ represent interactions: prey loss due to predation and predator growth from consuming prey. The downward arrow on $y$ represents natural mortality.
3.0.0.2 Generalized Logistic Model
The Generalized Logistic Model (GLM) is a flexible extension of the logistic function used to model S-shaped epidemic growth curves. It is given by
$\frac{dC}{dt}=r\,C^{p}(t)\left(1-\frac{C(t)}{K}\right),$ (3)
with initial condition
$C(0)=C_{0},$ (4)
where $C(t)$ is the cumulative number of cases at time $t$, $r$ is the generalized growth rate, $K$ is the final epidemic size, and $p\in[0,1]$ controls the growth dynamics: $p=0$ corresponds to constant incidence, $0<p<1$ to sub-exponential growth, and $p=1$ to exponential growth. Here, $C_{0}$ represents the first reported case. Moreover, $\frac{dC}{dt}$ serves as the observation operator mapping the model states to the reported incidence data. The GLM was chosen because it provides a parsimonious, single-equation description of epidemic trajectories and serves as a tractable case for testing inference accuracy and uncertainty when the full time series is directly observed. This model has also been widely applied in epidemiological settings to capture early and full-phase epidemic dynamics [85] . The GLM has been widely used as a phenomenological growth model to capture epidemic curves with rapid increase and saturation [85, 86] .
[…]
Figure 3: Diagram of the GLM. The circle represents the cumulative cases $C(t)$. The self-loop indicates growth governed by the generalized logistic equation. The dashed arrow indicates the observed incidence.
3.0.0.3 SEIUR Model
The SEIUR model tracks susceptible ($S$), exposed ($E$), reported infectious ($I$), unreported infectious ($U$), recovered ($R$), and cumulative reported cases ($C$). The dynamics are
$\begin{gathered}\frac{dS}{dt}=-\beta_{f}(t)\frac{(I+U)S}{N},\quad\frac{dE}{dt}=\beta_{f}(t)\frac{(I+U)S}{N}-\kappa E,\quad\frac{dI}{dt}=\kappa\rho E-\gamma I,\\ \frac{dU}{dt}=\kappa(1-\rho)E-\gamma U,\quad\frac{dR}{dt}=\gamma(I+U),\quad\frac{dC}{dt}=\kappa\rho E,\end{gathered}$ (5)
with initial conditions
$(S(0),E(0),I(0),U(0),R(0),C(0))=(N-C_{0},0,C_{0},0,0,C_{0}),$ (6)
where $N$ is the total population, $\beta_{f}(t)$ is the time-dependent transmission rate, $\kappa$ is the incubation rate, $\gamma$ is the recovery rate, $\rho$ is the reporting proportion, and $C_{0}$ is the first reported case. The initial condition assumes that all initially infected individuals are reported, and the remaining population is susceptible. The time-dependent transmission rate $\beta_{f}(t)$ is defined as the following
$\beta_{f}(t)=\begin{cases}\beta_{0},&\text{if }t<t_{\text{int}},\\[6.0pt] \beta_{1}+(\beta_{0}-\beta_{1})\,e^{-q_{1}(t-t_{\text{int}})},&\text{if }t\geq t_{\text{int}}.\end{cases}$
The time-dependent transmission rate $\beta_{f}(t)$ is modeled in a piecewise manner to reflect changes in population behavior and control measures following an intervention at time $t_{\mathrm{int}}$. Before the intervention, transmission is assumed to be constant at level $\beta_{0}$, representing relatively uncontrolled epidemic spread. After $t_{\mathrm{int}}$, transmission decreases smoothly toward a lower asymptotic level $\beta_{1}$ through an exponential decay, capturing the gradual adoption and increasing effectiveness of interventions such as mobility reduction, masking, or behavioral adaptation. The decay rate $q_{1}$ controls how rapidly these effects take hold, allowing the model to represent both abrupt and slow transitions while remaining parsimonious and identifiable.
This model represents a more realistic epidemic process with both observed and hidden states, making it ideal for testing how Bayesian and Frequentist methods handle latent variables, parameter coupling, and partial identifiability.
[…]
Figure 4: Compartmental diagram of the SEIUR model with underreporting. Circles represent the epidemiological compartments. Solid arrows indicate transitions between compartments, and the dashed arrow indicates the source of observed cases.
4 Methods
This section describes the Bayesian and Frequentist estimation frameworks used for parameter inference, uncertainty quantification, and forecasting. Let $Y=(y_{t_{1}},\ldots,y_{t_{n}})$ denote the observed data at discrete time points $t_{1},\ldots,t_{n}$, and let $\boldsymbol{\theta}$ represent the vector of unknown model parameters to be estimated. For the Lotka–Volterra (LV) model, $\boldsymbol{\theta}=(\alpha,\beta,\gamma,\delta)$. For the generalized logistic model (GLM), $\boldsymbol{\theta}=(r,p,K)$. For the SEIUR model, $\boldsymbol{\theta}=(\beta_{0},\beta_{1},q_{1},\rho,\kappa,\gamma_{1})$. Both estimation approaches assume a normal error structure to ensure comparability of results.
4.1 Bayesian Inference
Bayesian inference integrates prior knowledge with observed data to obtain the posterior distribution of model parameters [87]. This probabilistic framework is particularly effective when prior information is available or when data are sparse, noisy, or subject to partial observability, as it provides comprehensive uncertainty quantification through posterior distributions [88, 89].
4.1.1 Bayes’ Rule
According to Bayes’ theorem, the posterior distribution of parameters given the data is proportional to the product of the prior distribution and the likelihood:
$p(\boldsymbol{\theta}\mid Y)\propto p(\boldsymbol{\theta})\,p(Y\mid\boldsymbol{\theta}),$ (7)
where $p(\boldsymbol{\theta})$ is the prior distribution encoding existing knowledge about the parameters before observing data, $p(Y\mid\boldsymbol{\theta})$ is the likelihood function representing the probability of observing the data given parameter values, and $p(\boldsymbol{\theta}\mid Y)$ is the posterior distribution of parameters after incorporating the observed data.
4.1.2 Likelihood Function
We assume that observation errors are independent and identically distributed (i.i.d.) following a normal distribution with constant variance. For the Lotka–Volterra model with observed prey and predator populations, the likelihood is:
$y_{t_{j}}^{\text{prey}}\mid\boldsymbol{\theta}\sim\mathcal{N}(x_{\boldsymbol{\theta}}(t_{j}),\sigma_{\text{prey}}^{2}),\quad y_{t_{j}}^{\text{predator}}\mid\boldsymbol{\theta}\sim\mathcal{N}(y_{\boldsymbol{\theta}}(t_{j}),\sigma_{\text{predator}}^{2}),$ (8)
independently for $j=1,\ldots,n$, where $x_{\boldsymbol{\theta}}(t)$ and $y_{\boldsymbol{\theta}}(t)$ denote the prey and predator population solutions to the Lotka–Volterra ODE system given parameters $\boldsymbol{\theta}$, and $\sigma_{\text{prey}}^{2}$ and $\sigma_{\text{predator}}^{2}$ are observation error variances.
For the GLM, the likelihood is based on incident cases (the time derivative of cumulative cases):
$y_{t_{j}}\mid\boldsymbol{\theta}\sim\mathcal{N}\left(\frac{dC_{\boldsymbol{\theta}}(t_{j})}{dt},\sigma^{2}\right),\quad j=1,\ldots,n,$ (9)
where $C_{\boldsymbol{\theta}}(t)$ is the solution to the GLM ODE given parameters $\boldsymbol{\theta}$.
For the SEIUR model, the likelihood is based on incident reported cases:
$y_{t_{j}}\mid\boldsymbol{\theta}\sim\mathcal{N}\left(\kappa\rho E_{\boldsymbol{\theta}}(t_{j}),\sigma^{2}\right),\quad j=1,\ldots,n,$ (10)
where $E_{\boldsymbol{\theta}}(t)$ is the exposed compartment solution to the SEIUR ODE system given parameters $\boldsymbol{\theta}$, and $\kappa\rho E$ represents the flow of individuals from exposed to reported infectious status.
4.1.3 Prior Distributions
Prior distributions encode existing knowledge or beliefs about parameter values before observing data [44] . For parameters with established biological or epidemiological interpretations, we specify informative priors based on previous studies. When prior knowledge is limited, we use weakly informative or uniform priors over plausible parameter ranges. For structural identifiability analysis with simulated data, priors were centered on true parameter values with varying degrees of precision to assess the impact of prior strength on parameter recovery. Table 28 summarizes the prior distributions used for each model. The normal error variance $\sigma^{2}$ is assigned a weakly informative inverse-gamma or half-Cauchy prior to allow the data to inform the observation error scale.
4.1.4 Posterior Sampling and Convergence Diagnostics
Posterior distributions are approximated using Markov Chain Monte Carlo (MCMC) sampling implemented through the Hamiltonian Monte Carlo (HMC) algorithm in Stan [62, 65]. For each model, we ran four independent chains with a sufficient number of iterations (typically 2,000–4,000) after a burn-in period to ensure convergence to the stationary distribution. Convergence was assessed using the Gelman–Rubin diagnostic $\hat{R}$ statistic [65], which compares between-chain and within-chain variance; values of $\hat{R}<1.05$ (ideally $\hat{R}\approx 1.01$) indicate successful convergence. Effective sample size (ESS) was monitored to ensure adequate sampling of the posterior distribution. Trace plots and posterior density plots were visually inspected to confirm mixing and convergence.
Posterior summaries are reported as medians with 95% prediction intervals (PIs), which represent the central 95% of the posterior distribution. For forecasting, we propagate parameter uncertainty by sampling parameters values from the posterior distribution, solving the ODE for each sample, and summarizing the resulting forecast distribution with median and 95% prediction intervals (CIs).
4.1.5 Computational Implementation
Bayesian inference was conducted using the (BFF) toolbox [26], an R package designed for fitting and forecasting ODE-based epidemic models. The BFF toolbox provides an automated workflow that generates Stan code based on user-specified model structures, priors, and data inputs, eliminating the need for users to program directly in Stan. The toolbox outputs include posterior parameter distributions, convergence diagnostics ($\hat{R}$, ESS), trace plots, posterior density plots, and forecasting results with quantified uncertainty. Performance metrics including mean absolute error (MAE), mean squared error (MSE), weighted interval score (WIS), and 95% PI coverage are automatically computed to evaluate model fit and forecast accuracy.
4.2 Frequentist Inference
Frequentist estimation treats model parameters as fixed but unknown quantities and estimates them by optimizing an objective function based solely on the observed data, without incorporating prior information [33, 34] . Uncertainty quantification is performed through bootstrap resampling procedures.
4.2.1 Parameter Estimation
Under the assumption of normally distributed observation errors with constant variance (consistent with the Bayesian approach), parameter estimates are obtained by minimizing the sum of squared residuals between observed data and model predictions. This corresponds to nonlinear least squares (NLS) estimation:
$\hat{\boldsymbol{\theta}}=\underset{\boldsymbol{\theta}}{\text{argmin}}\sum_{j=1}^{n}\left(y_{t_{j}}-\mu_{j}(\boldsymbol{\theta})\right)^{2},$ (11)
where $\mu_{j}(\boldsymbol{\theta})$ is the model-predicted value at time $t_{j}$ given parameters $\boldsymbol{\theta}$. For the Lotka–Volterra model, $\mu_{j}(\boldsymbol{\theta})$ consists of predicted prey and predator populations $(x_{\boldsymbol{\theta}}(t_{j}),y_{\boldsymbol{\theta}}(t_{j}))$. For the GLM, $\mu_{j}(\boldsymbol{\theta})=dC_{\boldsymbol{\theta}}(t_{j})/dt$. For the SEIUR model, $\mu_{j}(\boldsymbol{\theta})=\kappa\rho E_{\boldsymbol{\theta}}(t_{j})$.
In our study, the Frequentist inference is performed using MATLAB’s fmincon with the Sequential Quadratic Programming (SQP) algorithm. To mitigate sensitivity to initial conditions and reduce the risk of convergence to local minima, the optimization is wrapped within a MultiStart framework using a CustomStartPointSet, with an optional RandomStartPointSet.
The observed data variance is estimated as:
$\hat{\sigma}^{2}=\frac{1}{n}\sum_{j=1}^{n}\left(y_{t_{j}}-\mu_{j}(\hat{\boldsymbol{\theta}})\right)^{2}.$ (12)
4.2.2 Uncertainty Quantification via Parametric Bootstrap
To quantify parameter uncertainty and construct prediction intervals, we employ a parametric bootstrap procedure [33, 35]. This approach generates synthetic datasets by resampling from the fitted model and re-estimates parameters for each synthetic dataset. The distribution of bootstrap parameter estimates characterizes sampling variability. The steps are as follows:
  1. 1.
    Generate bootstrap samples: For each bootstrap replicate $b=1,\ldots,B$ (typically B=200B=200500500), generate a synthetic dataset $\{y_{t_{1}}^{b},\ldots,y_{t_{n}}^{b}\}$ by sampling from the fitted model:
    $y_{t_{j}}^{b}\sim\mathcal{N}\left(\mu_{j}(\hat{\boldsymbol{\theta}}),\hat{\sigma}^{2}\right),\quad j=1,\ldots,n,$ (13)
    where $\hat{\boldsymbol{\theta}}$ and $\hat{\sigma}^{2}$ are the parameter estimates from the original data.
  2. 2.
    Re-estimate parameters: For each bootstrap sample $\{y_{t_{1}}^{b},\ldots,y_{t_{n}}^{b}\}$, solve the NLS optimization problem (Eq. 11) to obtain bootstrap parameter estimates $\hat{\boldsymbol{\theta}}^{b}$.
  3. 3.
    Construct prediction intervals: Use the empirical distribution of $\{\hat{\boldsymbol{\theta}}^{b}:b=1,\ldots,B\}$ to construct prediction intervals. For example, the 95% prediction interval for parameter $\theta_{k}$ is given by the 2.5th and 97.5th percentiles of $\{\hat{\theta}_{k}^{b}:b=1,\ldots,B\}$.
For forecasting, we propagate parameter uncertainty by solving the ODE model for each bootstrap parameter estimate $\hat{\boldsymbol{\theta}}^{b}$ to obtain forecast trajectories $\hat{y}^{b}(t_{n+h})$ for $h$-step-ahead predictions. The 95% prediction interval for $y(t_{n+h})$ is constructed from the 2.5th and 97.5th percentiles of $\{\hat{y}^{b}(t_{n+h}):b=1,\ldots,B\}$.
4.2.3 Computational Implementation
Frequentist inference was implemented using the QuantDiffForecast (QDF) MATLAB toolbox [25], which provides a comprehensive framework for parameter estimation, uncertainty quantification, and forecasting for ODE models. The QDF toolbox supports multiple optimization algorithms, flexible error structures (normal, Poisson, negative binomial), user-defined ODE systems, and automated bootstrap-based uncertainty quantification. The toolbox outputs parameter estimates with prediction intervals, fitted model trajectories, forecast distributions, and performance metrics (MAE, MSE, WIS, 95% PI coverage).
4.3 Performance Metrics
[…]
4.3.1 Mean Absolute Error (MAE)
MAE measures the average absolute deviation between model predictions and observed data:
$\text{MAE}=\frac{1}{N}\sum_{i=1}^{N}\left|f(t_{i},\hat{\boldsymbol{\theta}})-y_{t_{i}}\right|.$ (14)
MAE provides a direct measure of forecast accuracy on the original data scale. Lower MAE values indicate better point forecast performance. MAE is less sensitive to outliers compared to MSE.
4.3.2 Mean Squared Error (MSE)
MSE measures the average squared deviation between model predictions and observed data:
$\text{MSE}=\frac{1}{N}\sum_{i=1}^{N}\left(f(t_{i},\hat{\boldsymbol{\theta}})-y_{t_{i}}\right)^{2}.$ (15)
MSE penalizes larger errors more heavily than MAE due to the squaring operation, making it more sensitive to outliers. Lower MSE values indicate better fit. The square root of MSE (RMSE) is sometimes reported to return the metric to the original data scale.
4.3.3 Coverage of the 95% prediction interval
The 95% PI coverage quantifies the proportion of observed data points that fall within the 95% prediction interval, providing a measure of uncertainty calibration:
$\text{95\% PI Coverage}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(L_{t_{i}}<y_{t_{i}}<U_{t_{i}}),$ (16)
where $L_{t_{i}}$ and $U_{t_{i}}$ are the lower and upper bounds of the 95% prediction interval at time $t_{i}$, and $\mathbf{1}(\cdot)$ is an indicator function equal to 1 if the condition is true and 0 otherwise. Ideally, the coverage should be close to 95%. Coverage significantly below 95% indicates overly narrow intervals (underestimation of uncertainty), while coverage above 95% suggests overly wide intervals (overestimation of uncertainty).
4.3.4 Weighted Interval Score (WIS)
WIS is a proper scoring rule that evaluates the quality of the entire predictive distribution by combining sharpness (interval width) and calibration (penalties for observations outside intervals) [74, 73]. The interval score (IS) for a single prediction interval at level $\alpha$ is:
$\text{IS}_{\alpha}(F,y)=(u-l)+\frac{2}{\alpha}(l-y)\mathbf{1}(y<l)+\frac{2}{\alpha}(y-u)\mathbf{1}(y>u),$ (17)
where $l$ and $u$ represent the $\frac{\alpha}{2}$ and $\left(1-\frac{\alpha}{2}\right)$ quantiles of the forecast distribution $F$, respectively. The IS consists of three components:
  • Sharpness: (ul)(u-l), the width of the central $(1-\alpha)\times 100\%$ prediction interval. Narrower intervals receive lower scores, rewarding precise forecasts.
  • Underprediction penalty: $\frac{2}{\alpha}(l-y)\mathbf{1}(y<l)$, penalizes observations falling below the lower bound $l$, with the penalty proportional to the distance (ly)(l-y).
  • Overprediction penalty: $\frac{2}{\alpha}(y-u)\mathbf{1}(y>u)$, penalizes observations exceeding the upper bound $u$, with the penalty proportional to the distance (yu)(y-u).
To comprehensively evaluate the full predictive distribution, we compute the WIS by averaging interval scores over multiple interval levels $(1-\alpha_{1})<(1-\alpha_{2})<\cdots<(1-\alpha_{K})$ along with the predictive median $\tilde{y}$ (which can be viewed as an interval at level $(1-\alpha_{0})\to 0$):
$\text{WIS}_{\alpha_{0:K}}(F,y)=\frac{1}{K+\frac{1}{2}}\left(w_{0}|y-\tilde{y}|+\sum_{k=1}^{K}w_{k}\,\text{IS}_{\alpha_{k}}(F,y)\right),$ (18)
where $w_{k}=\frac{\alpha_{k}}{2}$ for $k=1,\ldots,K$ and $w_{0}=\frac{1}{2}$. WIS quantifies how close the entire predictive distribution is to the observed data in units on the original data scale. Lower WIS values indicate better forecast performance, balancing sharpness and calibration.
4.3.5 Interpretation
Lower values of MAE, MSE, and WIS indicate better model performance. For 95% PI coverage, values close to 95% indicate well-calibrated uncertainty estimates. Together, these metrics provide a comprehensive assessment of both point forecast accuracy (MAE, MSE), distributional forecast quality (WIS), and uncertainty calibration (95% PI coverage). We report these metrics separately for the calibration period (in-sample fit) and forecasting period (out-of-sample prediction) to distinguish between model fitting and predictive performance.
5 Structural Identifiability
Structural identifiability (SI) analysis determines whether model parameters can, in principle, be uniquely recovered from perfect, noise-free observations of the system outputs, given the model structure and observation scheme [27, 71]. A parameter is structurally identifiable if its value can be uniquely determined from the input-output relationship defined by the model equations and the set of observable variables. Conversely, a parameter is structurally unidentifiable if multiple distinct parameter values produce identical model outputs, making unique parameter recovery impossible regardless of data quality or quantity.
Structural identifiability is a prerequisite for meaningful parameter estimation: if a parameter is structurally unidentifiable, no amount of data or sophisticated inference methods can recover its true value [28, 27] . SI analysis is particularly important in compartmental models where only a subset of state variables is observed, as partial observability often leads to identifiability deficits [72] . Understanding which parameters are identifiable under different observation scenarios guides experimental design, informs prior specification in Bayesian inference, and helps interpret estimation results.
5.1 Methodology
We conducted symbolic structural identifiability analysis using StructuralIdentifiability.jl [71], a Julia package that employs differential algebra methods to determine identifiability of parameters in ODE models. For each model, we analyzed two scenarios: (1) known initial conditions (ICs), where the initial values of all state variables are assumed to be known exactly, and (2) unknown initial conditions, where ICs are treated as additional unknown parameters to be estimated. The distinction between known and unknown ICs is critical, as identifiability results can differ substantially depending on whether ICs must be inferred from data [70].
For the (LV) model, we examined three observation schemes to assess how partial observability affects identifiability:
  • LV-1 (Both prey and predator observed): Both $x(t)$ (prey) and $y(t)$ (predator) populations are observed.
  • LV-2 (Predator only): Only $y(t)$ (predator) is observed; prey population $x(t)$ is unobserved.
  • LV-3 (Prey only): Only $x(t)$ (prey) is observed; predator population $y(t)$ is unobserved.
For the Generalized Logistic Model (GLM), we analyzed the identifiability when observing incident cases (the time derivative of cumulative cases $C(t)$). For the SEIUR model, we considered the scenario where only incident reported cases (the flow $\kappa\rho E(t)$) are observed, representing typical epidemic surveillance data where only a fraction of infections are detected and reported.
5.2 Results
Table 2 summarizes the structural identifiability results for all models under different observation schemes and initial condition assumptions.
Table 2: Structural identifiability results for all models under different observation schemes. The table indicates which parameters and state variables are identifiable or unidentifiable when initial conditions (ICs) are known versus unknown. LV-1, LV-2, and LV-3 refer to Lotka–Volterra scenarios with both species observed, predator only, and prey only, respectively. “NULL” indicates that no states or parameters are unidentifiable.
Model Observations Identifiable (unknown IC) Unidentifiable (unknown IC) Identifiable (known IC) Unidentifiable (known IC)
LV-1 $x(t),y(t)$ $\alpha,\beta,\delta,\gamma,x(t),y(t)$ All states and parameters
LV-2 $y(t)$ only $\alpha,\beta,\gamma,y(t)$ $\delta,x(t)$ All states and parameters
LV-3 $x(t)$ only $\alpha,\delta,\gamma,x(t)$ $\beta,y(t)$ All states and parameters
GLM C(t),dC/dtC(t),\,dC/dt C,r,K,pC,\,r,\,K,\,p All states and parameters
SEIUR $\kappa\rho E(t)$ $C,\,\kappa,\,\gamma_{1}$ $S,\,E,\,I,\,U,\,R,\,N,\,\rho$ All states and parameters
5.3 Interpretation
5.3.0.1 Lotka–Volterra Model.
When both predator and prey time series are observed (LV-1), all model parameters ($\alpha,\beta,\gamma,\delta$) are structurally identifiable regardless of whether initial conditions are known or unknown. The coupled dynamics of the two species provide sufficient information to uniquely determine all interaction parameters.
With partial observation, identifiability losses emerge when initial conditions are unknown. When only the predator is observed (LV-2), the prey growth-from-consumption parameter $\delta$ and the prey population trajectory $x(t)$ become unidentifiable. The observed predator dynamics constrain only the product $\delta x(t)$ rather than $\delta$ and $x(t)$ separately, creating a structural non-uniqueness. Similarly, when only the prey is observed (LV-3), the predation rate $\beta$ and the predator population $y(t)$ are unidentifiable because the prey dynamics constrain only the product $\beta y(t)$.
Importantly, when initial conditions are known, all parameters become identifiable even under partial observation scenarios (LV-2 and LV-3). Knowledge of the initial prey and predator populations breaks the symmetry and allows unique recovery of all parameters. This highlights the value of accurate initial condition estimates in ecological field studies.
5.3.0.2 Generalized Logistic Model.
For the GLM, observation of cumulative cases $C(t)$ and incident cases $dC/dt$ yields full structural identifiability of all parameters $(r,p,K)$ under both known and unknown initial conditions. The single-equation structure of the GLM, combined with direct observation of both the state variable and its derivative, ensures that all growth parameters are uniquely determined. This robust identifiability makes the GLM a reliable choice for phenomenological epidemic modeling when the full case trajectory is observed.
5.3.0.3 SEIUR Model.
The SEIUR model exhibits substantial identifiability challenges due to its multi-compartment structure and partial observability. Even when all compartments $(S,E,I,U,R,C)$ are theoretically observable, only the cumulative reported cases $C$, the incubation rate $\kappa$, and the recovery rate $\gamma_{1}$ remain structurally identifiable when initial conditions are unknown. The remaining state variables ($S,E,I,U,R$) and critical parameters including the total population size $N$ and the reporting proportion $\rho$ are structurally unidentifiable.
This identifiability deficit arises because the observed incidence data (new reported cases per day) depend on the product $\kappa\rho E(t)$ rather than on $\kappa$, $\rho$, and $E(t)$ individually. Multiple combinations of reporting proportion $\rho$, exposed population $E(t)$, and total population $N$ can produce identical observed case counts. The unobserved unreported infectious compartment $U(t)$ further compounds the identifiability problem, as the split between reported and unreported infections is not directly constrained by the data.
When initial conditions are known, all SEIUR parameters become identifiable. This underscores the importance of accurate estimates of the initial susceptible and exposed populations (e.g., from seroprevalence surveys or contact tracing data) for reliable parameter inference in epidemic models with underreporting.
5.4 Implications for Inference
The structural identifiability analysis provides critical context for interpreting the empirical parameter estimates and forecasting results presented in the Result section. For the Lotka–Volterra model, we expect both Bayesian and Frequentist methods to successfully recover all parameters when both species are observed (LV-1), but parameter estimates may be unreliable or exhibit high uncertainty in the partial observation scenarios (LV-2, LV-3) where structural identifiability is compromised.
For the GLM applied to lung injury and mpox data, the robust structural identifiability suggests that both inference methods should yield well-constrained parameter estimates and reliable forecasts, provided the data quality is sufficient.
For the SEIUR model applied to COVID-19 data in Spain, the limited structural identifiability indicates that several model parameters cannot be uniquely determined from case incidence data alone. We expect parameter estimates for $\rho$, $N$, and the latent compartments to exhibit substantial uncertainty and potential non-uniqueness. In this setting, Bayesian inference may have an advantage by incorporating prior information to regularize the otherwise ill-posed inverse problem, while Frequentist methods may struggle without additional constraints or data sources.
These predictions will be assessed empirically in the Results section, where we compare parameter estimation accuracy, uncertainty quantification, and forecast performance across models and observation scenarios.
Summary of Models, Domains, Data Sources, and Observation Scenarios
As summarized in Table 3, the models, application domains, data sources, and observation scenarios used in this study are presented.
Table 3: The models, application domains, data sources, and observation scenarios used in this study. The datasets span ecological, clinical, and epidemiological systems.
Model Domain Data Source Scenarios
Lotka–Volterra Ecology Hudson–Bay lynx–hare population (predator + prey), Prey-only, Predator-only
GLM Clinical Lung injury progression Full data
GLM Epidemiology 2022 U.S. mpox outbreak Full data
SEIUR Epidemiology COVID-19 incidence in Spain First-wave data
6 Results
This section reports parameter estimates, calibration performance metrics, fitting visualization, and Bayesian convergence diagnostics across the models. We organize results by model and, when applicable, by observation scenario.
6.1 Lotka–Volterra Model
We fitted the Lotka–Volterra model to Hudson Bay lynx-hare data under three observation scenarios: (1) both prey and predator observed (LV-1), (2) predator only (LV-2), and (3) prey only (LV-3). These scenarios allow us to assess how partial observability affects parameter recovery and forecasting performance in light of the structural identifiability results presented in Section 5.
6.1.1 LV-1: Joint predator–prey observations
As shown in Figure 5, with both time series available, both methods recover the characteristic coupled oscillations with well-calibrated uncertainty. However, BFF provides better coverage of prey data between 1905-1910. Parameter estimates are almost similar for $\beta$ (0.02 vs. 0.025) and $\delta$ (both 0.03) but differ for other parameters, as shown in Table 4.
The calibration results show a trade-off that depends on the series (Table 5). For the predator series, QDF has lower errors, such as lower MAE and MSE. For the prey series, BFF performs better, with lower errors and higher coverage (100% compared to 90.48%). Table 6 also shows that the Bayesian chains mixed well, with $\hat{R}$ values close to 1 for all parameters.
Table 4: Parameter estimates for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing both predator and prey, using QDF and BFF approaches.
$\alpha$ $\beta$ $\delta$ $\gamma$
BFF 0.40 (0.36, 0.44) 0.02 (0.02, 0.02) 0.03 (0.03, 0.03) 1.06 (0.95, 1.17)
QDF 0.55 (0.51, 0.60) 0.025 (0.022, 0.029) 0.03 (0.02, 0.03) 0.84 (0.76, 0.92)
Figure 5: Fitting visualization for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing both predator and prey, using QDF and BFF approaches. Top row: BFF. Bottom row: QDF. Left column: predator. Right column: prey.
Table 5: Performance metrics for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing both predator and prey, using QDF and BFF approaches.
MAEMAE MSEMSE WISWIS $95\%PI$
Predator BFF 4.95 35.93 3.07 100
QDF 2.87 13.56 1.88 100
Prey BFF 2.24 8.04 1.46 100
QDF 3.29 21.96 2.27 90.48
Table 6: Convergence diagnostic for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing both predator and prey in Bayesian approach.
$\alpha$ $\beta$ $\delta$ $\gamma$
$\hat{R}$ 1.01 1.01 1.01 1.01
6.1.2 LV-2: Predator-only observations
We next fit the Lotka–Volterra model using predator observations only. As expected under reduced observability (Section 5), parameter uncertainty increases and multiple parameter combinations can explain the observed predator cycles (Table 7). Despite this, point-fit performance is similar across methods (Table 8): QDF has slightly lower MAE (3.8 vs. 3.6 for BFF), while BFF has slightly lower MSE and WIS. PI coverage remains high for both (100% for BFF; 95.24% for QDF), and the Bayesian chains show $\hat{R}=1$ in this fit (Table 9).
Table 7: Parameter estimates for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing only predator, using QDF and BFF approaches.
$\alpha$ $\beta$ $\delta$ $\gamma$
Predator BFF 1.20 (0.75, 1.90) 0.03 (0.01, 0.06) 0.02 (0.01, 0.03) 0.36 (0.22, 0.62)
QDF 0.97 (0.64, 1.00) 0.02 (0.02, 0.06) 0.025 (0.02, 0.036) 0.46 (0.40, 0.70)
BFF and QDF yield comparable predator-only fits (Figure 6), but implied prey dynamics are necessarily weakly constrained without prey observations.
Figure 6: Fitting visualization for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing only predator, using QDF and BFF approaches. Left BFF, and Right QDF.
Table 8: Performance metrics for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing only predator, using QDF and BFF approaches.
MAEMAE MSEMSE WISWIS $95\%PI$
Predator BFF 3.6 18.07 2.22 100
QDF 3.8 19.7 2.33 95.24
Table 9: Convergence diagnostic for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing only predator in Bayesian approach.
$\alpha$ $\beta$ $\delta$ $\gamma$
Predator $\hat{R}$ 1 1 1 1
6.1.3 LV-3: Prey-only observations
Finally, we fit the Lotka–Volterra model using prey observations only. In this scenario, QDF achieves lower point-error metrics (Table 11: MAE 3.3 vs. 4.6; MSE 15.83 vs. 27.34; WIS 2.1 vs. 2.8), while BFF attains higher PI coverage (100% vs. 90.48%). Parameter estimates remain plausible but are less constrained than in the joint case (Table 10), and the Bayesian fit again shows $\hat{R}\approx 1.01$ (Table 12).
Table 10: Parameter estimates for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing only prey, using QDF and BFF approaches.
$\alpha$ $\beta$ $\delta$ $\gamma$
Prey BFF 0.34 (0.28, 0.52) 0.01 (0.00, 0.05) 0.04 (0.02, 0.05) 1.40 (0.75, 1.90)
QDF 0.42 (0.34, 0.55) 0.013 (0.005, 0.035) 0.03 (0.02, 0.04) 1.20 (0.78, 1.50)
Both methods track the main prey fluctuations under prey-only calibration (Figure 7), but implied predator dynamics remain uncertain without direct predator observations.
Figure 7: Fitting visualization for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing only prey, using QDF and BFF approaches. Left BFF, and Right QDF.
MAEMAE MSEMSE WISWIS $95\%PI$
Prey BFF 4.6 27.34 2.8 100
QDF 3.3 15.83 2.1 90.48
Table 11: Performance metrics for the Lotka–Volterra model, with Hudson Bay lynx-hare data, observing only prey, using QDF and BFF approaches.
Table 12 displays the Gelman-Rubin diagnostic ($\hat{R}$) values for BFF, all of which are 1.01, indicating successful MCMC convergence.
Table 12: Convergence diagnostic for the Lotka–Volterra model, observing only prey in Bayesian approach.
$\alpha$ $\beta$ $\delta$ $\gamma$
Prey $\hat{R}$ 1.01 1.01 1.01 1.01
Across the three observation scenarios, parameter estimates and interval widths vary in a parameter-specific manner (Figure 8). Joint predator–prey observation primarily affects interaction-related parameters, most notably $\beta$, for which point estimates differ substantially across inference methods, reflecting stronger parameter coupling and improved identifiability. By contrast, for other parameters ($\alpha$, $\delta$, and $\gamma$), interval widths and estimates overlap considerably across observation scenarios, indicating that joint observation does not uniformly reduce uncertainty relative to single-series calibration.
Figure 8: Parameter comparison for the Lotka–Volterra model, with Hudson Bay lynx-hare data, across observation schemes: joint predator-prey, only predator, only prey. We use a log scale for the x-axis.
Comparing predator-only vs. joint data availability, adding prey observations improves calibration performance for both methods across MAE/MSE/WIS and PI coverage (Figure 9), underscoring the value of joint observations for dynamic ecological systems.
Figure 9: Predictive accuracy vs. data availability in LV. Across MAE, MSE, WIS, and 95% PI, joint predator–prey observation improves performance for both frameworks compared with predator-only data.
Likewise, adding predator observations to a prey-only fit improves performance across metrics for both frameworks (Figure 10).
Figure 10: Effect of adding predator–prey data when starting from prey-only series. Using both species lowers MAE/MSE, reduces WIS, and raises 95% PI for BFF and QDF relative to prey-only fits.
The histogram summaries (Figures 2025) reinforce these patterns: joint data produce tighter, more concentrated uncertainty summaries, while predator-only and prey-only settings yield broader distributions, consistent with reduced information under partial observation (Section 5).
Table 13: Prior distributions and bounds for the parameters in the Lotka–Volterra model. The priors are truncated normal distributions with mean 0.5 and standard deviation 0.5, constrained to be non-negative (T[0,]).
Parameter Prior Distribution Lower Bound (LB) Upper Bound (UB)
$\alpha$ normal(0.5, 0.5) T[0,] 0 NA
$\beta$ normal(0.5, 0.5) T[0,] 0 NA
$\delta$ normal(0.5, 0.5) T[0,] 0 NA
$\gamma$ normal(0.5, 0.5) T[0,] 0 NA
Table 14: Parameter ranges for the Lotka–Volterra model of (QDF).
Parameter Lower Bound (LB) Upper Bound (UB)
$\alpha$ 0.001 1.0
$\beta$ 0.001 0.5
$\delta$ 0.001 0.5
$\gamma$ 0.001 1.5
6.2 GLM model for Lung Injury
The GLM provides a well-identified baseline example in which the data support stable parameter recovery under both frameworks (Section 5). Both approaches capture the rise and decline in the lung-injury time series with QDF providing a better fit (Figure 11).
Figure 11: Fitting visualization for the GLM model, with lung-injury data, using QDF and BFF approaches. Left QDF, and Right BFF.
The inferred growth parameters are close across methods (Table 15), with $r$ near 0.32–0.33 and $p$ near 1.0, indicating consistent recovery of the growth shape. The main discrepancy is in the estimated carrying capacity $k$, where BFF yields a higher central estimate than QDF.
In calibration performance, QDF achieves lower point-error metrics (MAE 9.42 vs. 11.85; MSE 152.80 vs. 412.29; WIS 5.75 vs. 7.15) while both methods attain high PI coverage (Table 16). This pattern is consistent with an identifiable, low-dimensional setting where optimization-based estimation can yield very sharp point fits without compromising interval calibration.
The Bayesian fit converged well for this dataset ($\hat{R}=1$ for all GLM parameters; Table 17), supporting reliability of the reported posterior summaries.
Table 15: Parameter estimates for the GLM model, with lung-injury data, using QDF and BFF approaches.
r p k
BFF 0.33 (0.3,0.36) 0.98 (0.96,1) 2734.03 (2732.03,2735.82)
QDF 0.32 (0.30,0.35) 1 (0.99,1.02) 2549.5 (2438.9,2662.9)
Table 16: Performance metrics for the GLM model, with lung-injury data, using QDF and BFF approaches.
[…]
BFF 11.85 412.2943 7.15 97.83
QDF 9.42 152.80 5.75 100
Table 17: Convergence diagnostics for the GLM model, with lung-injury data, using QDF and BFF approaches.
r p k
$\hat{R}$ 1 1 1
The cross-method parameter comparison reinforces that $(r,p)$ are tightly constrained in both frameworks, with only modest differences in uncertainty widths (Figure 12).
Figure 12: Parameter comparison for the GLM model, with lung-injury data, using QDF and BFF approaches.
Across the summary metrics, the same ranking is apparent: QDF yields lower MAE/MSE/WIS while PI coverage remains comparable (Figure 13).
Figure 13: Performance metrics comparison for the GLM model, with lung-injury data, using QDF and BFF approaches.
The parameter bounds used for applying the BFF and QDF for this model is provided in Table 18 and 19.
Table 18: Parameter bounds for the GLM model, with lung-injury data, using QDF and BFF approaches.
Parameter Lower Bound (LB) Upper Bound (UB)
$r$ 0.0 NA
pp 0.0 NA
$k$ 0.0 NA
Table 19: Parameter bounds for the GLM model, with lung-injury data, using QDF and BFF approaches.
Parameter Lower Bound (LB) Upper Bound (UB)
$r$ 0.01 5.0
pp 0.0 2.0
$k$ 0.01 1000000.0
6.3 GLM model for mpox epidemic in the USA
For the 2022 U.S. mpox epidemic, both methods reproduce the overall rise–fall pattern, indicating that the GLM captures the main outbreak trajectory under either inference framework (Figure 14). Parameter estimates are broadly consistent across methods, with modest differences in uncertainty widths (Table 20). For example, $p$ is tightly concentrated around $\sim 0.8$–0.85 for both methods, while $r$ differs more in its central estimate (BFF 1.95 vs. QDF 2.25), reflecting method-specific tradeoffs in fitting the growth and saturation phases.
Figure 14: Fitting visualization for the GLM model, with U.S. mpox data, using QDF and BFF approaches. Left QDF, and Right BFF.
Calibration-period metrics favor QDF on point-error measures (MAE 9.45 vs. 11.739; MSE 150.47 vs. 399.43; WIS 5.8 vs. 7.175), while both methods attain high PI coverage (100% for QDF vs. 97.8% for BFF; Table 21). The Bayesian fit converged cleanly ($\hat{R}=1$ for all GLM parameters; Table 22), supporting stability of the reported BFF intervals.
Table 20: Parameter estimates for the GLM model, with U.S. mpox data, using QDF and BFF approaches.
r p K
BFF 1.95 (1.76,2.15) 0.84 (0.82,0.85) 29041.01 (27784.85,30164.68)
QDF 2.25 (2.01,2.7) 0.81 (0.80,0.83) 30363.6 (29168.42,31838.14)
The cross-method parameter comparison (Figure 15) and metric summary (Figure 16) reinforce these trends: parameter uncertainty is generally tight under both methods, and QDF attains lower MAE/MSE/WIS while maintaining near-nominal PI coverage.
Table 21: Performance metrics for the GLM model, with U.S. mpox data, using QDF and BFF approaches.
[…]
BFF 11.739 399.43 7.175 97.8
QDF 9.45 150.47 5.8 100
Table 22: Convergence diagnostics for the GLM model, with U.S. mpox data, using BFF approache.
r p k
$\hat{R}$ 1 1 1
Figure 15: Parameter comparison for the GLM model, with U.S. mpox data, using QDF and BFF approaches. We use a log scale for the x-axis.
Figure 16: Performance metrics comparison for the GLM model, with U.S. mpox data, using QDF and BFF approaches.
The parameter bounds used for applying the BFF and QDF for this model is provided in Table 23 and 24.
Table 23: Parameter ranges for the GLM on mpox data under BFF with Normal error.
Parameter Lower Bound (LB) Upper Bound (UB)
$r$ 0.0 NA
pp 0.0 NA
$k$ 0.0 30000
Table 24: Parameter ranges for the GLM on mpox data under QDF with Normal error.
Parameter Lower Bound (LB) Upper Bound (UB)
$r$ 0.01 5.0
pp 0.0 2.0
$k$ 0.01 1000000.0
6.4 SEIUR model of COVID-19 cases in Spain during the first wave
For Spain’s first COVID-19 wave, both methods track the rise and decline in reported incidence (Figure 17), but the inferred uncertainties and error metrics differ in ways consistent with the SEIUR model’s limited identifiability under incidence-only observations (Section 5). Compared with QDF, BFF yields lower errors (MAE 294.57 vs. 352.30; MSE 210,754 vs. 311,054; WIS 211.30 vs. 250.46) with similar PI coverage (94% vs. 92%; Table 26), suggesting an advantage for Bayesian regularization in this partially observed, higher-dimensional setting.
Figure 17: Fitting visualization for the SEIUR model, with Spain’s first COVID-19 wave data, using QDF and BFF approaches. Left QDF, and Right BFF.
The parameter summaries (Table 25) differ most for $\rho$ and $q_{1}$, which are strongly coupled to unobserved compartments and therefore weakly constrained by incidence alone. This helps explain why the two frameworks can yield different central estimates despite producing similar fitted trajectories.
Convergence diagnostics indicate some remaining sampling difficulty for the Bayesian SEIUR fit (Table 27; several parameters have $\hat{R}>1.1$), which is expected in a coupled, weakly identified setting. Accordingly, we interpret SEIUR posterior summaries with appropriate caution.
$\beta_{0}$ $\beta_{1}$ $q_{1}$ $\rho$ $\kappa$ $\gamma_{1}$
BFF 2.18 (1.6, 3.09) 1.69 (1.14, 2.24) 0.88 (0.64, 1.41) 0.6 (0.28, 0.92) 2.39 (1.23, 3.73) 1.78 (1.22, 2.59)
QDF 1.93 (1.18, 1.96) 1.35 (0.54, 1.38) 1.99 (1.13, 2.00) 0.96 (0.72, 1.00) 1.62 (0.61, 1.75) 1.44 (0.65, 1.47)
Table 25: Parameter estimates for the SEIUR model, with Spain’s first COVID-19 wave data, using QDF and BFF approaches.
Table 26: Performance metrics for the SEIUR model, with Spain’s first COVID-19 wave data, using QDF and BFF approaches.
[…]
BFF 294.57 210754.1222 211.303 94
QDF 352.2978 311053.5723 250.4568 92
Table 27: Convergence diagnostics for the SEIUR model, with Spain’s first COVID-19 wave data, using BFF approach.
$\beta_{0}$ $\beta_{1}$ $q_{1}$ $\rho$ $\kappa$ $\gamma_{1}$
$\hat{R}$ 1.18 1.22 1.17 1 1.07 1.21
The parameter comparison (Figure 18) highlights broader uncertainty for several rates under BFF, consistent with posterior uncertainty propagation under identifiability limitations.
Figure 18: Parameter comparison for the SEIUR model, with Spain’s first COVID-19 wave data, using BFF approach.
The metric summary (Figure 19) reinforces the table-based ranking: BFF achieves lower MAE/MSE/WIS while PI coverage remains close to nominal for both methods.
Figure 19: Performance metrics comparison for the SEIUR model, with Spain’s first COVID-19 wave data, using BFF approach.
The parameter bounds used for applying the BFF and QDF for this model is provided in Table 28 and 29.
Table 28: Parameter ranges for the SEIUR model, with Spain’s first COVID-19 wave data, using BFF approach.
Parameter Lower Bound (LB) Upper Bound (UB)
$\beta_{0}$ 0.0 NA
$\beta_{1}$ 0.0 NA
$q1$ 0.0 NA
$\rho$ 0.0 NA
$\kappa$ 0.0 NA
$\gamma_{1}$ 0.0 NA
NN 0.0 47000000
Table 29: Parameter ranges for the SEIUR model, with Spain’s first COVID-19 wave data, using QDF approach.
Parameter Lower Bound (LB) Upper Bound (UB)
$\beta_{0}$ 0.001 2.0
$\beta_{1}$ 0.001 2.0
$q1$ 0.001 2.0
$\rho$ 0.001 1.0
$\kappa$ 0.001 2.0
$\gamma_{1}$ 0.001 4.0
NN 50000 1000000000.0
7 Discussion
This study compares Bayesian and Frequentist inference across ecological, clinical, and epidemiological models, with a focus on how data availability, model structure, and identifiability shape inference performance. Across all examples, we find that neither framework universally outperforms the other. Instead, performance depends critically on the degree of structural identifiability, the presence of latent states, and whether the primary goal is accurate point prediction or reliable uncertainty quantification. Bayesian methods, implemented via BayesianFitForecast (BFF), consistently provide stronger uncertainty estimates and diagnostic tools, while Frequentist methods, implemented via QuantDiffForecast (QDF), often achieve lower point prediction errors with substantially lower computational cost when identifiability is strong.
For the Lotka–Volterra model, inference performance varied markedly across observation scenarios. When both prey and predator populations were observed, both methods produced stable and plausible parameter estimates. QDF yielded slightly lower point errors for predator dynamics, while BFF achieved better uncertainty calibration for prey. When only a single population was observed, QDF continued to provide sharper point predictions, whereas BFF maintained superior prediction interval coverage. Bayesian convergence diagnostics indicated stable posterior sampling even under partial observation, confirming the reliability of Bayesian inference in these settings.
Differences in parameter estimates across observation schemes reflect changes in practical identifiability. With full observation of both species, interaction and turnover parameters are more strongly informed by the data, reducing confounding but also exposing differences between inference philosophies. In this regime, Bayesian and optimization-based approaches may favor different trade-offs when fitting coupled dynamics, leading to larger discrepancies in some parameters. Under partial observation, the Lotka–Volterra system becomes more weakly identifiable, as multiple parameter combinations can reproduce similar marginal dynamics for the observed series while uncertainty is absorbed by the unobserved state. In this case, both methods are effectively constrained toward similar parameter regimes through regularization, bounds, or prior structure, which can make several parameter estimates appear closer despite increased uncertainty.
These empirical findings closely follow the predictions of the structural identifiability analysis. When both prey and predator are observed, all Lotka–Volterra parameters are structurally identifiable, whereas observing only one species leads to non-identifiability unless initial conditions are known. This limitation affects both Bayesian and Frequentist approaches equally, underscoring that observability places fundamental limits on inference quality regardless of statistical framework.
For the generalized logistic model applied to lung injury and mpox outbreaks, QDF consistently achieved better point forecast accuracy while maintaining similar or better uncertainty calibration. In the mpox case, QDF produced lower MAE, MSE, and WIS, while achieving near-perfect or perfect prediction interval coverage. Bayesian inference showed excellent convergence for all parameters, confirming well-sampled posteriors. Similar patterns were observed for lung injury. These results are explained by the strong structural identifiability of the generalized logistic model: the single-equation structure and direct observation of incident cases ensure that all parameters are identifiable, even when initial conditions are unknown. In this setting, optimization-based inference provides accurate estimates efficiently, while Bayesian inference adds computational overhead without clear gains in predictive accuracy.
A contrasting pattern emerges for the SEIUR model applied to COVID-19 data from Spain. Here, BFF outperformed QDF across all metrics, including point accuracy and uncertainty calibration. The SEIUR model includes latent compartments and exhibits limited structural identifiability when only reported cases are observed and initial conditions are unknown. Structural identifiability analysis shows that most parameters and state variables are unidentifiable in this setting, with only a small subset remaining identifiable. Under these conditions, Bayesian inference benefits from prior information, which regularizes the ill-posed inverse problem and stabilizes both parameter estimates and forecasts. In contrast, deterministic optimization must search a high-dimensional and weakly constrained parameter space, leading to less reliable results.
Taken together, these results show that structural identifiability sets hard limits on what can be learned from data. No statistical method can recover parameters that are structurally unidentifiable, and identifiability analysis should therefore precede model fitting. When identifiability is strong and data are rich, Frequentist methods are well suited for minimizing point forecast error and computational cost. When identifiability is weak, data are sparse or noisy, or models include latent states, Bayesian methods provide more reliable uncertainty quantification and stable inference.
Several methodological choices also influence these results. The use of different software environments reflects methodological requirements rather than implementation convenience: BFF relies on Stan-based Bayesian inference supported in R, while QDF depends on deterministic optimization and repeated numerical ODE solutions, for which MATLAB provides highly optimized solvers. Both approaches use identical model structures, data inputs, and evaluation metrics, ensuring that performance differences arise from inference methodology rather than computational platform.
In the practical identifiability analysis, initial conditions were treated as known and fixed to align with the structural identifiability setting and to isolate differences arising from inference methodology. Allowing initial conditions to be estimated would introduce additional confounding, particularly in partially observed systems, and would be expected to increase uncertainty and variability in parameter estimates. Both QDF and BFF support joint estimation of initial conditions, and examining the impact of this added uncertainty remains an important direction for future work.
This study intentionally focuses on low-dimensional models to allow transparent interpretation and the application of differential-algebraic structural identifiability methods. While this enables a clean methodological comparison, it also imposes limitations. We assumed normal error models, analyzed systems independently, and considered idealized identifiability conditions. For larger and more realistic models, practical identifiability tools based on Fisher information or sensitivity analysis may be necessary, especially under noisy or incomplete observations. Similarly, while Hamiltonian Monte Carlo performs well for moderate parameter dimensions, higher-dimensional models may require alternative sampling strategies such as adaptive Metropolis or DRAM.
Future work could explore hybrid inference approaches that combine the computational efficiency of optimization with Bayesian uncertainty quantification, as well as extensions to higher-dimensional, spatial, stochastic, or agent-based models. Structural identifiability analysis could also be integrated more directly into experimental design, guiding data collection before inference is attempted. Incorporating machine learning approaches such as neural ordinary differential equations or physics-informed neural networks may provide complementary advantages when combined with identifiability-aware modeling.
Supplementary Material
7.1 LV-Model (BFF) Histograms
Figure 20: Histograms of the Lotka–Volterra parameters $(\alpha,\beta,\delta,\gamma)$.
Figure 21: Histograms for predator data only of $(\alpha,\beta,\delta,\gamma)$ Parameters
Figure 22: Histograms for prey data only of $(\alpha,\beta,\delta,\gamma)$ Parameters
7.2 LV-model (QDF) Histograms
Figure 23: Histograms for predator-prey data of $(\alpha,\beta,\delta,\gamma)$ Parameters
Figure 24: Histograms for predator data only of $(\alpha,\beta,\delta,\gamma)$ Parameters
[…]
7.3 GLM-Model-Lung Injury (BFF) Histograms
Figure 26: Histograms for GLM data of r, k, and p Parameters
7.4 GLM-Model-Lung Injury (QDF) Histograms
Figure 27: Histograms r,p, and k Parameters
7.5 GLM-Model-Mpox (BFF) Histograms
Figure 28: Histograms for GLM data of r, k, and p Parameters
7.6 GLM-Model-Mpox (QDF) Histograms
Figure 29: Histograms for GLM data of r, k, and p Parameters
7.7 SEIUR-Model (BFF) Histograms
Figure 30: Histograms ($\beta_{0}$, $\beta_{1}$, $q$, rhorho, kappakappa, and $sigma_{1}$ ) Parameters
7.8 SEIUR-Model (QDF) Histograms
Figure 31: Histograms for ($\beta_{0}$, $\beta_{1}$, $q$, $\rho$, $\kappa$, and $\gamma_{1}$ ) Parameters